### v1

# Separable Operator Networks

Abstract
Operator learning has become a powerful tool in machine learning for modeling complex physical systems. Although Deep Operator Networks (DeepONet) show promise, they require extensive data acquisition. Physics-informed DeepONets (PI-DeepONet) mitigate data scarcity but suffer from inefficient training processes. We introduce Separable Operator Networks (SepONet), a novel framework that significantly enhances the efficiency of physics-informed operator learning. SepONet uses independent trunk networks to learn basis functions separately for different coordinate axes, enabling faster and more memory-efficient training via forward-mode automatic differentiation. We provide theoretical guarantees for SepONet using the universal approximation theorem and validate its performance through comprehensive benchmarking against PI-DeepONet. Our results demonstrate that for the 1D time-dependent advection equation, when targeting a mean relative $\ell_{2}$ error of less than 6% on 100 unseen variable coefficients, SepONet provides up to 112× training speed-up and 82× GPU memory usage reduction compared to PI-DeepONet. Similar computational advantages are observed across various partial differential equations, with SepONet’s efficiency gains scaling favorably as problem complexity increases. This work paves the way for extreme-scale learning of continuous mappings between infinite-dimensional function spaces.
Keywords Partial differential equations  $\cdot$ Separation of variables  $\cdot$ Scientific machine learning  $\cdot$ Operator learning
1 Introduction
Operator learning, which aims to learn mappings between infinite-dimensional function spaces, has gained significant attention in scientific machine learning thanks to its ability to model complex dynamics in physics systems. This approach has been successfully applied to a wide range of applications, including climate modeling [1, 2], multiphysics simulation [3, 4, 5, 6, 7], inverse design [8, 9] and more [10, 11, 12, 13]. Various operator learning algorithms [14, 15, 16, 17, 18, 19, 20, 21, 22] have been developed to tackle these applications, with Deep Operator Networks (DeepONets) [14] being particularly notable due to their universal approximation guarantee for operators [23] and robustness [24].
To approximate the function operator $G:\mathcal{U}\rightarrow\mathcal{S}$, a DeepONet consists of three key components: an encoder mapping the input function $u\in\mathcal{U}$ to its point-wise evaluations at fixed sensor locations $x_{1},x_{2},\ldots,x_{m}$, a branch net processing these evaluations to produce $r$ coefficients $b_{1},b_{2},\ldots,b_{r}$, and a trunk net taking continuous coordinate $y\in\boldsymbol{R}^{d}$ to return $r$ basis functions $t_{1}(y),t_{2}(y),\ldots,t_{r}(y)$. The output function $s\in\mathcal{S}$ is approximated as $s(u)(y)\approx\sum_{k=1}^{r}b_{k}(u(x_{1}),u(x_{2}),\ldots,u(x_{m}))t_{k}(y)$. In the context of parametric partial differential equations (PDEs), $u$ represents the PDE configuration and $s$ represents the corresponding solution. For example, in thermal simulations [3], $u$ might represent heat sources and boundary conditions that parameterize the heat equation, while $s$ is the resulting temperature field solution.
DeepONets are usually trained in a supervised manner using a dataset of $N_{f}$ pair functions $\left\{u^{(i)},s^{(i)}\right\}_{i=1}^{N_{f}}$. Unlike traditional numerical methods which require repeated simulations for each different PDE configuration, once well trained, a DeepONet allows for efficient parallel inference in abstract infinite-dimensional function spaces. Given new PDE configurations, it can immediately provide the corresponding solutions. However, this advantage comes with a significant challenge: to achieve satisfactory generalization error, the number of required input-output function pairs grows exponentially [14, 25, 26]. Generating enough function pairs can be computationally expensive or even impractical in some applications, creating a bottleneck in the effective deployment of DeepONets.
Physics-informed deep operator networks (PI-DeepONet) [27] have been introduced as a solution to the costly data acquisition problem. Inspired by physics-informed neural networks (PINNs) [28], PI-DeepONet learns operators by constraining the DeepONet outputs to approximately satisfy the underlying governing PDE system parameterized by the input function $u$. This is achieved by penalizing the physics loss (including PDE residual loss, initial loss and boundary loss), thus eliminating the need for ground-truth output functions $s$. However, PI-DeepONet shares the same disadvantage as PINNs in that the training process is both memory-intensive and time-consuming [29, 30]. This inefficiency, common to both PI-DeepONet and PINNs, arises from the need to compute high-order derivatives of the PDE solution with respect to numerous collocation points during physics loss optimization. This computation typically relies on reverse-mode automatic differentiation [31], involving the backpropagation of physics loss through the unrolled computational graph to update model parameters. The inefficiency is even more pronounced for PI-DeepONet, as it requires evaluating the physics loss across multiple PDE configurations. While numerous studies [32, 33, 29, 34, 35, 30, 36] have proposed methods to enhance PINN training efficiency, there has been limited research focused on improving the training efficiency of PI-DeepONet specifically.
To address the inefficiency in training PI-DeepONet, we propose Separable Operator Networks (SepONet), inspired by the separation of variables technique in solving PDEs and recent work on separable PINN [30]. SepONet employs $d$ independent trunk nets to learn basis functions separately for different axes. In this framework, a SepONet prediction of the output function given input function $u$ is given by $s(u)(y)\approx\sum_{k=1}^{r}b_{k}(u(x_{1}),u(x_{2}),\ldots,u(x_{m}))\prod_{n=1}^{d}t_{k}^{n}(y_{n})$ for $y=(y_{1},y_{2},\ldots,y_{d})$. This simple yet effective modification not only facilitates efficient learning of variables with potentially different scales, but also enables fast training and memory-efficient implementation of SepONet by leveraging forward-mode automatic differentiation [37]. Our key contributions are:
  1. 1.
    We introduce SepONet, a physics-informed operator learning framework that significantly enhances scaling efficiency in terms of training time and GPU memory usage, enabling extreme-scale learning of continuous mappings between infinite-dimensional function spaces.
  2. 2.
    We provide a theoretical foundation for SepONet through the universal approximation theorem, demonstrating its capability to approximate any nonlinear continuous operator with arbitrary accuracy.
  3. 3.
    We validate our theoretical results through benchmarking SepONet against PI-DeepONet on a range of 1D time-dependent PDEs. Our findings reveal that scaling up SepONet consistently improves its accuracy and typically outperforms PI-DeepONet, achieving up to 112× training speed-up with minimal memory increase for these problems. Notably, at moderately large scales where PI-DeepONet exhausts 80GB of GPU memory, SepONet operates efficiently with less than 1GB.
2 Preliminaries
2.1 Operator Learning for Solving Parametric Partial Differential Equations
Let $\mathcal{X}$ and $\mathcal{Y}$ be Banach spaces, with $K\subseteq\mathcal{X}$ and $K_{1}\subseteq\mathcal{Y}$ being compact sets. Consider a nonlinear continuous operator $G:\mathcal{U}\rightarrow\mathcal{S}$, mapping functions from one infinite-dimensional space to another, where $\mathcal{U}\subseteq C(K)$ and $\mathcal{S}\subseteq C(K_{1})$. The goal of operator learning is to approximate the operator $G$ using neural networks parameterized by $\boldsymbol{\theta}$, denoted as $G_{\boldsymbol{\theta}}$. Here, $\mathcal{U}$ and $\mathcal{S}$ represent spaces of functions where the input and output functions have dimensions $d_{u}$ and $d_{s}$, respectively. We focus on the scalar case where $d_{u}=d_{s}=1$ throughout this paper without loss of generality; however, it should be noted that the results apply to arbitrary $d_{u}$ and $d_{s}$. A widely used framework for approximating such an operator $G$ involves constructing $G_{\boldsymbol{\theta}}$ through three maps [38]:
$G_{\boldsymbol{\theta}}\approx G:=\mathcal{D}\circ\mathcal{A}\circ\mathcal{E}.$ (1)
First, the encoder $\mathcal{E}:\mathcal{U}\rightarrow\boldsymbol{R}^{m}$ maps an input function $u\in\mathcal{U}$ to a finite-dimensional feature representation. Next, the approximator $\mathcal{A}:\boldsymbol{R}^{m}\rightarrow\boldsymbol{R}^{r}$ transforms this encoded data within the finite-dimensional space $\boldsymbol{R}^{m}$ to another finite-dimensional space $\boldsymbol{R}^{r}$. Finally, the decoder $\mathcal{D}:\boldsymbol{R}^{r}\rightarrow\mathcal{S}$ produces the output function $s(y)=G(u)(y)$ for $y\in K_{1}$.
In the context of solving parametric partial differential equations (PDEs), consider PDEs of the form:
$\mathcal{N}(u,s)$ $=0,\;\mathcal{I}(u,s)=0,\;\mathcal{B}(u,s)=0,$ (2)
where $\mathcal{N}$ is a nonlinear differential operator, $\mathcal{I}$ and $\mathcal{B}$ represent the initial and boundary conditions, $u\in\mathcal{U}$ denotes the PDE configurations, and $s\in\mathcal{S}$ denotes the corresponding PDE solution. Assuming that for any $u\in\mathcal{U}$ there exists a unique solution $s(u)\in\mathcal{S}$, we can define the solution operator $G:\mathcal{U}\rightarrow\mathcal{S}$ as:
$G(u)=s(u).$ (3)
The goal is to learn the operator $G$ that maps PDE configuration $u(x)$ for $x\in K$ (e.g., source terms, PDE coefficients, initial conditions) to the PDE solution $s(y)$. Various operator learning algorithms [14, 16, 39, 17, 18, 19, 20, 21, 22] have been developed to approximate the solution operator defined in (3). Among these, Deep Operator Networks (DeepONet) [14] are particularly notable due to their universal approximation capability of approximating operators [23], robustness [24], and the foundation they provide for numerous other approaches [40, 41, 42, 43, 44]. As such, we focus on the DeepONet framework, which we will introduce in the following section.
2.2 Deep Operator Networks (DeepONet)
The original DeepONet formulation [14] can be analyzed through the 3-step approximation framework 1. The encoder maps the input function $u$ to its point-wise evaluations at $m$ sensors $x_{1},x_{2},\ldots,x_{m}\in K$. Two separate neural networks (usually multilayer perceptrons), the branch net and the trunk net, serve as the approximator and decoder, respectively. The branch net processes $\left(u(x_{1}),u(x_{2}),\ldots,u(x_{m})\right)$ to produce a feature embedding $\left(b_{1},b_{2},\ldots,b_{r}\right)$. The trunk net, taking the continuous coordinate $y$ as input, outputs a feature embedding $\left(t_{1},t_{2},\ldots,t_{r}\right)$. The final DeepONet prediction of a function $u$ for a query $y$ is:
$G_{\boldsymbol{\theta}}(u)(y)=\sum_{k=1}^{r}b_{k}(u(x_{1}),u(x_{2}),\ldots,u(x_{m}))t_{k}(y),$ (4)
where $\boldsymbol{\theta}$ represents all the trainable parameters in the branch and trunk nets. Given training data consisting of input functions $\left\{u^{(i)}\right\}_{i=1}^{N_{f}}$ sampled from $\mathcal{U}$ and corresponding output functions $\left\{s(u^{(i)})\right\}_{i=1}^{N_{f}}$ evaluated at collocation points $\left\{y^{(j)}\right\}_{j=1}^{N_{c}}$, the parameters $\boldsymbol{\theta}$ can be optimized using gradient descent by minimizing the operator loss:
$\mathcal{L}_{operator}=\frac{1}{N_{f}N_{c}}\sum_{i=1}^{N_{f}}\sum_{j=1}^{N_{c}}\Big{|}G_{\boldsymbol{\theta}}(u^{(i)})(y^{(j)})-s(u^{(i)})(y^{(j)})\Big{|}^{2}.$ (5)
Once trained, a DeepONet allows for efficient inference in abstract infinite-dimensional function spaces. Given new PDE configurations, it can immediately provide the corresponding solutions, in contrast to traditional numerical methods, which require repeated simulations for each different PDE configuration.
2.3 Physics-Informed Deep Operator Networks (PI-DeepONet)
Despite DeepONet’s remarkable success across a range of applications in multiphysics simulation [4, 5, 6], inverse design [8], and carbon storage [10], its training process is highly dependent on the availability of training data, which can be very costly. On one hand, the generalization error of DeepONets scales with the number of training input-output function pairs [14, 25, 26]. On the other hand, generating a large number of high-quality training data is expensive or even impractical in some applications. For example, in simulating high Reynolds number ($Re$) turbulent flow [45], accurate numerical simulations require a fine mesh, leading to a computational cost scaling with $Re^{3}$ [46], making the generation of sufficiently large and diverse training datasets prohibitively expensive.
To address the need for costly data acquisition, physics-informed deep operator networks (PI-DeepONet) [27], inspired by physics-informed neural networks (PINNs) [28], have been proposed to learn operators without relying on observed data. PI-DeepONets are governed by the PDE and its corresponding initial and boundary conditions. They are trained by minimizing the following physics loss:
$\mathcal{L}_{physics}=\mathcal{L}_{residual}+\lambda_{I}\mathcal{L}_{initial}+\lambda_{b}\mathcal{L}_{boundary},$ (6)
where each loss term is defined based on the underlying PDE:
$\mathcal{L}_{residual}$ $=\frac{1}{N_{f}N_{r}}\sum_{i=1}^{N_{f}}\sum_{j=1}^{N_{r}}\Big{|}\mathcal{N}(u^{(i)},G_{\boldsymbol{\theta}}(u^{(i)})(y_{r}^{(j)}))\Big{|}^{2},$ (7)
$\mathcal{L}_{initial}$ $=\frac{1}{N_{f}N_{I}}\sum_{i=1}^{N_{f}}\sum_{j=1}^{N_{I}}\Big{|}\mathcal{I}(u^{(i)},G_{\boldsymbol{\theta}}(u^{(i)})(y_{I}^{(j)}))\Big{|}^{2},$
$\mathcal{L}_{boundary}$ $=\frac{1}{N_{f}N_{b}}\sum_{i=1}^{N_{f}}\sum_{j=1}^{N_{b}}\Big{|}\mathcal{B}(u^{(i)},G_{\boldsymbol{\theta}}(u^{(i)})(y_{b}^{(j)}))\Big{|}^{2}.$
Here, $\left\{u^{(i)}\right\}_{i=1}^{N_{f}}$ are the input training functions sampled from $\mathcal{U}$. $\bigl{\{}y_{r}^{(j)}\bigl{\}}_{j=1}^{N_{r}}$, $\bigl{\{}y_{I}^{(j)}\bigl{\}}_{j=1}^{N_{I}}$, and $\bigl{\{}y_{b}^{(j)}\bigl{\}}_{j=1}^{N_{b}}$ represent the residual points, initial points, and boundary points, respectively. $\lambda_{I}$ and $\lambda_{b}$ denote the weight coefficients for different loss terms. An example of the training points for PI-DeepONet is illustrated in Fig. 1.
Figure 1: Training points for PI-DeepONet, including randomly sampled initial, boundary, and residual points.
3 Inefficiency in Training PI-DeepONet
Given a parametric PDE system (2), PI-DeepONet can be trained to learn its solution operator $G:\mathcal{U}\rightarrow\mathcal{S}$ in an unsupervised manner. However, as noted in the original PI-DeepONet paper [27], the training process can be both memory-intensive and time-consuming. Similar to PINNs [28], this inefficiency arises because optimizing the physics loss requires calculating high-order derivatives of the PDE solution with respect to numerous collocation points, typically achieved via reverse-mode automatic differentiation [31]. This process involves backpropagating the physics loss through the unrolled computational graph to update the model parameters. For PI-DeepONet, the inefficiency is even more pronounced, as the physics loss (Eq. (7)) must be evaluated across multiple PDE configurations. Although various works [32, 29, 30] have proposed different methods to improve the training efficiency of PINNs, little research has focused on enhancing the training efficiency of PI-DeepONet.
3.1 A Motivating Example
To investigate how the performance and training cost of PI-DeepONet scale with the training data size, let us consider a simple 1D linear heat equation:
$\frac{\partial s}{\partial t}=\frac{1}{\pi^{2}}\frac{\partial^{2}s}{\partial x^{2}},\quad(x,t)\in(0,1)\times(0,1],$ (8)
$s(x,0)=u(x),\quad x\in(0,1),$
$s(0,t)=s(1,t)=0,\quad t\in(0,1).$
The goal is to learn the operator from the initial condition $u(x)$ to the solution $s(x,t)$. We approximate the solution operator by a PI-DeepONet $G_{\boldsymbol{\theta}}$, where the encoder maps the input function $u$ to its point-wise evaluations at 128 equi-spaced sensors in $[0,1]$. The branch and trunk nets are two separate 5-layer fully-connected neural networks with 50 neurons per hidden layer. The output dimensions of both networks are 50 (rank $r=50$). We apply Tanh activation for the branch net and Sine activation for the trunk net. The PI-DeepONet is trained by minimizing the physics loss (Eq. (6)) using gradient descent with the Adam optimizer [47] for 80,000 iterations. The initial learning rate is $1\times 10^{-3}$ and decays by a factor 0.9 every 1,000 iterations. Specifically, we set $\lambda_{I}=20,\lambda_{b}=1$ in (6). The loss terms in Eq. (7) are defined as:
$\mathcal{L}_{residual}$ $=\frac{1}{N_{f}N_{r}}\sum_{i=1}^{N_{f}}\sum_{j=1}^{N_{r}}\bigg{|}\frac{\partial G_{\boldsymbol{\theta}}(u^{(i)})(x_{r}^{(j)},t_{r}^{(j)})}{\partial t_{r}^{(j)}}-\frac{1}{\pi^{2}}\frac{\partial^{2}G_{\boldsymbol{\theta}}(u^{(i)})(x_{r}^{(j)},t_{r}^{(j)})}{\partial\bigl{(}x_{r}^{(j)}\bigl{)}^{2}}\bigg{|}^{2},$ (9)
$\mathcal{L}_{initial}$ $=\frac{1}{N_{f}N_{I}}\sum_{i=1}^{N_{f}}\sum_{j=1}^{N_{I}}\Big{|}G_{\boldsymbol{\theta}}(u^{(i)})(x_{I}^{(j)},0)-u^{(i)}(x_{I}^{(j)})\Big{|}^{2},$
$\mathcal{L}_{boundary}$ $=\frac{1}{N_{f}N_{b}}\sum_{i=1}^{N_{f}}\sum_{j=1}^{N_{b}}\Bigl{(}\Big{|}G_{\boldsymbol{\theta}}(u^{(i)})(0,t_{b}^{(j)})\Big{|}^{2}+\Big{|}G_{\boldsymbol{\theta}}(u^{(i)})(1,t_{b}^{(j)})\Big{|}^{2}\Bigl{)}.$
Here the training functions (initial conditions) $\left\{u^{(i)}\right\}_{i=1}^{N_{f}}$ are generated from a Gaussian random field (GRF) ~ $\mathcal{N}\left(0,25^{2}\left(-\Delta+5^{2}I\right)^{-4}\right)$ using the Chebfun package [48], satisfying the zero Dirichlet boundary condition. For every input training function $u^{(i)}$, the sets $\bigl{\{}(x_{r}^{(j)},t_{r}^{(j)})\bigl{\}}_{j=1}^{N_{r}}$, $\bigl{\{}(x_{I}^{(j)},0)\bigl{\}}_{j=1}^{N_{I}}$, and $\bigl{\{}(0,t_{b}^{(j)}),(1,t_{b}^{(j)})\bigl{\}}_{j=1}^{N_{b}}$ are randomly sampled within the computation domain to enforce the PDE residual, initial and boundary conditions, respectively. To balance the interplay of different terms in the loss function, we set $N_{r}=N_{I}^{2}=N_{b}^{2}=N_{c}$, where $N_{c}$ is referred to as the number of training points. We also resample input training functions and training points (including residual, initial and boundary points) every 100 iterations. The code is implemented using JAX and Equinox libraries [49, 50] and all training is performed on a single NVIDIA A100 GPU with 80 GB memory.
To evaluate the generalization error and training cost of PI-DeepONet, we sample 100 unseen initial conditions from the same GRF and apply Runge-Kutta forward Euler method to obtain the corresponding solutions on a $128\times 128$ uniform spatio-temporal grid. Fixing the number of input training functions at $N_{f}=100$ and varying the number of training points $N_{c}$, the results are reported in Table 1. The mean and standard deviation of the relative $\ell_{2}$ error and root mean squared error (RMSE) are calculated over 100 initial conditions in the test dataset. As $N_{c}$ increases, test accuracy improves, but GPU memory usage and training time increase dramatically.
Similarly, as shown in Table 2, when $N_{c}=128^{2}$ is fixed and $N_{f}$ varies, the same trend is observed. These results demonstrate that while PI-DeepONet is effective, it can become extremely memory-intensive and time-consuming as the number of training points and training functions increases to enhance accuracy.
Table 1: Performance scaling of PI-DeepONet on the heat equation with varying number of training points ($N_{c}$) and fixed number of input training functions ($N_{f}=100$).
Performance Metrics \ $N_{c}$ $8^{2}$ $16^{2}$ $32^{2}$ $64^{2}$ $128^{2}$
Relative $\ell_{2}$ error (%) $1.93\pm 1.81$ $1.61\pm 1.35$ $1.47\pm 1.00$ $1.23\pm 0.73$ $0.94\pm 0.40$
RMSE (×$\times 10^{-2}$) $0.33\pm 0.15$ $0.27\pm 0.13$ $0.25\pm 0.12$ $0.19\pm 0.09$ $0.15\pm 0.07$
Memory (GB) 0.711 1.225 3.023 9.207 35.563
Training time (hours) 0.0601 0.0883 0.2181 1.0804 4.1531
Table 2: Performance scaling of PI-DeepONet on the heat equation with varying number of input training functions ($N_{f}$) and fixed number of training points ($N_{c}=128^{2}$).
Performance Metrics \ $N_{f}$ $5$ $10$ $20$ $50$ $100$
Relative $\ell_{2}$ error (%) $2.64\pm 2.13$ $2.03\pm 1.97$ $1.51\pm 0.96$ $1.39\pm 0.87$ $0.94\pm 0.40$
RMSE (×$\times 10^{-2}$) $0.60\pm 0.63$ $0.35\pm 0.18$ $0.34\pm 0.19$ $0.30\pm 0.17$ $0.15\pm 0.07$
Memory (GB) 2.765 5.071 9.173 17.919 35.563
Training time (hours) 0.1816 0.3224 0.8664 2.0721 4.1531
3.2 Solving the Motivation Example by Separation of Variables
The method of separation of variables has been widely used in solving linear PDEs. Consider the heat equation (8) presented in section 3.1, we seek a solution in the form:
$s(x,t)=X(x)T(t)$ (10)
for functions $X$, $T$ to be determined. Substituting (10) into (8) yields:
$\frac{X^{{}^{\prime\prime}}}{X}=-\lambda~{}~{}\text{and}~{}~{}\frac{\pi^{2}T^{{}^{\prime}}}{T}=-\lambda$ (11)
for some constant $\lambda$. To satisfy the boundary condition, $X$ must solve the following eigenvalue problem:
$X^{{}^{\prime\prime}}(x)+\lambda X(x)$ $=0,\quad x\in(0,1),$ (12)
$X(0)=X(1)$ $=0,$
and $T$ must solve the ODE problem:
$T^{{}^{\prime}}(t)=-\frac{\lambda}{\pi^{2}}T(t).$ (13)
The eigenvalue problem (12) has a sequence of solutions:
$\lambda_{k}=(k\pi)^{2},\quad X_{k}(x)=\sin(k\pi x),\quad\text{for}~{}~{}k=1,2,\dots$ (14)
For any $\lambda$, the ODE solution for $T$ is $T(t)=Ae^{-\frac{\lambda}{\pi^{2}}t}$ for some constant $A$. Thus, for each eigenfunction $X_{k}$ with corresponding eigenvalue $\lambda_{k}$, we have a solution $T_{k}$ such that the function
$s_{k}(x,t)=X_{k}(x)T_{k}(t)$ (15)
will be a solution of (11). In fact, an infinite series of the form
$s(x,t)=\sum_{k=1}^{\infty}X_{k}(x)T_{k}(t)=\sum_{k=1}^{\infty}A_{k}e^{-k^{2}t}\sin(k\pi x)$ (16)
will also be a solution satisfying the differential operator and boundary condition of the heat equation (8) subject to appropriate convergence assumptions of this series. Now let $s(x,0)=u(x)$, we can find coefficients:
[…]
such that (16) is the exact solution of the heat equation (8).
Bridging Separation of Variables Techniques with Operator Learning.
Given some initial condition $u$, recall that a (PI-)DeepONet predicts its corresponding solution for a query $y=(x,t)$ as:
$G_{\boldsymbol{\theta}}(u)(x,t)=\sum_{k=1}^{r}b_{k}(u(x_{1}),u(x_{2}),\ldots,u(x_{m}))t_{k}(x,t).$ (18)
This formulation actually resembles the structure of the solution Eq. (16) obtained via separation of variables techniques. In an ideal case, a (PI-)DeepONet will exactly mimic the solution operator if the branch net learns the integral operator and the trunk net learns the basis functions of $x$ and $t$, that is:
$b_{k}(u(x_{1}),u(x_{2}),\ldots,u(x_{m}))\equiv A_{k},\quad t_{k}(x,t)\equiv e^{-k^{2}t}\sin(k\pi x),\quad\text{for}~{}~{}k=1,2,\dots$ (19)
As discussed in Section 3.1, achieving accurate approximations of the integral operator and basis functions (19) in practice requires a greater number of input training functions and training points. While this improves the model’s ability to generalize to unseen initial conditions, it significantly increases training costs. This poses a substantial challenge to the practical implementation of operator learning in large-scale applications such as 3D integrated circuits design [26] and computational fluid dynamics simulation [46]. Therefore, developing a fast-to-train, memory-efficient, and accurate operator learning model is crucial.
4 Separable Operator Networks (SepONet)
4.1 Key Idea
Inspired by the method of separation of variables and recent work on separable PINN [30], we propose using separable operator networks (SepONet) to learn basis functions separately for different axes. For instance, a SepONet approximation to the solution operator of the heat equation (8) is given by:
$G_{\boldsymbol{\theta}}(u)(x,t)=\sum_{k=1}^{r}b_{k}(u(x_{1}),u(x_{2}),\ldots,u(x_{m}))\tau_{k}(t)\zeta_{k}(x).$ (20)
Here, $b_{k}$ is the $k$-th output of the branch net, as in DeepONet. However, unlike DeepONet, which employs a single trunk net that processes both $t$ and $x$ together, SepONet uses two independent trunk nets. The functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ are the $k$-th outputs of these trunk nets, taking the temporal input $t$ and spatial input $x$ separately. More generally, if the target output function $G(u)(y)$ is a multivariate function with $d$ variables such that $y=(y_{1},y_{2},\ldots,y_{d})\in\boldsymbol{R}^{d}$, a SepONet approximation for the operator $G:\mathcal{U}\rightarrow\mathcal{S}$ takes the form:
$G_{\boldsymbol{\theta}}(u)(y)=\sum_{k=1}^{r}\underbrace{b_{k}(u(x_{1}),u(x_{2}),\ldots,u(x_{m}))}_{branch}\prod_{n=1}^{d}\underbrace{t_{k}^{n}(y_{n})}_{trunk_{n}},$ (21)
where $t_{k}^{n}(y_{n})$ denotes the $k$-th output of the $n$-th trunk net.
SepONet can also be analyzed through the 3-step approximation framework (1). The encoder and approximator are identical to those in DeepONet. The decoder in SepONet consists of $d$ fully-connected neural networks (trunk nets), each learning $r$ basis functions for the corresponding variable $y_{n}$. For each $k$, the $d$ basis functions $t_{k}^{n}(y_{n})$ are multiplied together. These products are then linearly combined using the coefficients $b_{k}$, which are the outputs of the branch net, to obtain the final prediction as shown in Eq. 21.
This formulation can be viewed as a finite-dimensional parameterization of the solution operator by truncating the basis function series at a maximal number of ranks $r$. SepONet not only enjoys the advantage that basis functions for different variables with potentially different scales can be learned more easily and efficiently, but also allows for fast and efficient training by leveraging forward-mode automatic differentiation, which we will discuss in the following sections. It is worth noting that if one is only interested in solving deterministic PDEs under a certain configuration (i.e., $u$ is fixed), then the coefficients $b_{k}(u)$ are also fixed and can be absorbed by the basis functions. In this case, SepONet will reduce to separable PINN [30], which have been proven to be efficient and accurate in solving single-scenario PDE systems [51, 52].
4.2 SepONet Architecture and Implementation Details
For simplicity, we use the heat equation example (8) to present the network architecture of SepONet and illustrate how this simple modification to PI-DeepONet allows for fast training and memory-efficient implementation. Suppose we have $N_{f}$ input functions and need to evaluate the SepONet $G_{\boldsymbol{\theta}}$ for each input function on the same $N^{2}$ residual points $\bigl{\{}(x_{r}^{(j)},t_{r}^{(j)})\bigl{\}}_{j=1}^{N^{2}}$, $N$ boundary points $\bigl{\{}(0,t_{b}^{(j)}),(1,t_{b}^{(j)})\bigl{\}}_{j=1}^{N}$ on each boundary, and $N$ initial points $\bigl{\{}(x_{I}^{(j)},0)\bigl{\}}_{j=1}^{N}$ to calculate each loss term defined in Eq. (9).
Data Sampling.
Given some input function $u$, to evaluate the output function $G_{\boldsymbol{\theta}}(u)$ on $N^{2}$ residual points, SepONet employs a different sampling strategy than PI-DeepONet. Instead of directly sampling $N^{2}$ points from the computation domain $(0,1)\times(0,1]$ (as shown in Fig. 1), SepONet randomly samples $N$ points $\bigl{\{}x_{r}^{(i)}\bigl{\}}_{i=1}^{N}$ from $(0,1)$ and $\bigl{\{}t_{r}^{(i)}\bigl{\}}_{i=1}^{N}$ from $(0,1]$ separately (as shown in the left panel of Fig. 2). The initial points $\bigl{\{}(x_{I}^{(j)},0)\bigl{\}}_{j=1}^{N}$ and boundary points $\bigl{\{}(0,t_{b}^{(j)}),(1,t_{b}^{(j)})\bigl{\}}_{j=1}^{N}$ are randomly sampled from the computation domain, similar to PI-DeepONet.
Forward Pass.
The forward pass of SepONet, illustrated in Fig. 3, follows the formulation (21). It begins with the encoder $\mathcal{E}$ mapping the input function $u$ to its point-wise evaluations at $m$ sensor locations, producing evaluations $u_{1},u_{2},\ldots,u_{m}$. These evaluations are processed by the branch net $\mathcal{A}$, which outputs $r$ coefficients $b_{1},b_{2},\ldots,b_{r}$.
Next, the trunk nets process the temporal inputs $t^{(1)},t^{(2)},\ldots,t^{(N)}$ and spatial inputs $x^{(1)},x^{(2)},\ldots,x^{(N)}$ separately. Specifically, Trunk Net 1 outputs evaluations $\tau_{k}(t^{(1)}),\tau_{k}(t^{(2)}),\ldots,\tau_{k}(t^{(N)})$ for each of the $r$ basis functions $k=1,2,\ldots,r$, resulting in an $r\times N$ tensor $\boldsymbol{\tau}=(\tau_{1}^{(:)},\tau_{2}^{(:)},\ldots,\tau_{r}^{(:)})^{\top}$. Similarly, Trunk Net 2 outputs evaluations $\zeta_{k}(x^{(1)}),\zeta_{k}(x^{(2)}),\ldots,\zeta_{k}(x^{(N)})$ for each of the $r$ basis functions $k=1,2,\ldots,r$, resulting in another $r\times N$ tensor $\boldsymbol{\zeta}=(\zeta_{1}^{(:)},\zeta_{2}^{(:)},\ldots,\zeta_{r}^{(:)})^{\top}$.
These outputs are combined using an outer product $\bigotimes$ to form separable basis functions for each mode $k$, resulting in a tensor of size $r\times N\times N$, where the $(k,i,j)$-th entry is given by $\tau_{k}(t^{(i)})\cdot\zeta_{k}(x^{(j)})$. This tensor represents the evaluations of the basis functions on the mesh grid created by the $t$ and $x$ inputs, as shown in the right panel of Figure 2.
The decoder $\mathcal{D}$ then linearly combines these separable basis functions using the coefficients from the branch net to obtain the final predictions. This is efficiently achieved using the Einstein summation convention (einsum), which combines the $r$ coefficients $\boldsymbol{b}=(b_{1},b_{2},\ldots,b_{r})$ with the $r\times N\times N$ tensor to produce an output tensor of shape $N\times N$. The $(i,j)$-th entry of this tensor represents the approximated output function $G_{\boldsymbol{\theta}}(u)$ evaluated at $x=x^{(j)}$, $t=t^{(i)}$.
For $N_{f}$ input functions, the same einsum procedure is applied. The $N_{f}\times r$ coefficients are combined with the $r\times N\times N$ tensor to produce an output tensor of shape $N_{f}\times N\times N$. The $(k,i,j)$-th entry of this tensor represents the predicted output function $G_{\boldsymbol{\theta}}(u^{(k)})$ for the input function $u^{(k)}$, evaluated at $x=x^{(j)}$, $t=t^{(i)}$.
Backpropagation.
To evaluate the physics loss (Eq. (9)), SepONet’s output tensor, with shape $N_{f}\times N\times N$, enables more efficient computation of high-order derivatives using forward-mode automatic differentiation (AD) [37] compared to PI-DeepONet’s $N_{f}\times N^{2}\times 1$ output. Forward-mode AD is particularly advantageous for large numbers of functions and points, as it computes derivatives along each axis independently. In contrast, PI-DeepONet requires substantial computational overhead to compute high-order derivatives with respect to each collocation point in the flattened structure via reverse-mode AD, resulting in significant memory and time consumption. For a more detailed explanation of forward- and reverse-mode AD, we refer readers to [30, 53]. Once the physics loss is computed, reverse-mode AD is employed to update the model parameters.
Inference.
Once trained, SepONet can efficiently solve the heat equation (8) for any number of new initial conditions on any size mesh grid. This is achieved by combining the learned coefficients from the branch net with the basis functions learned by the trunk nets. An example of the learned basis functions for the heat equation is shown in Fig. 4. In this instance, we set $r=50$ and present a subset of the basis functions. While some basis functions for $x$ and $t$ (e.g., $\zeta_{46}(x)$ and $\tau_{46}(t)$) resemble the expected sine and exponential functions from Eq. (16), the SepONet basis functions do not necessarily converge to the analytic solution obtained by separation of variables in general. A more detailed discussion about the SepONet basis functions can be found in Appendix B.
Figure 2: Left: SepONet’s residual points sampling by randomly selecting $N$ points for each axis. Right: Resulting mesh grid of residual points for SepONet. The initial and boundary points for SepONet are also randomly sampled, similar to PI-DeepONet. The mesh grid in the right panel represents where the output function is evaluated, although the actual input points for SepONet are only $N$ points on each axis.
Figure 3: Network architecture of SepONet. The encoder $\mathcal{E}$ maps the input function $u$ to $m$ sensor evaluations $u_{1},\ldots,u_{m}$. The branch net $\mathcal{A}$ outputs $r$ coefficients $b_{1},\ldots,b_{r}$. Trunk Net 1 processes temporal inputs $t^{(1)},\ldots,t^{(N)}$ and Trunk Net 2 processes spatial inputs $x^{(1)},\ldots,x^{(N)}$, producing tensors $\boldsymbol{\tau}$ and $\boldsymbol{\zeta}$ of shape $r\times N$. These are combined using an outer product $\bigotimes$, resulting in a tensor of shape $r\times N\times N$. The decoder $\mathcal{D}$ then uses Einstein summation (einsum) to combine these basis functions with the branch net coefficients, producing an output tensor of shape $N\times N$. For $N_{f}$ input functions, the output tensor shape is $N_{f}\times N\times N$.
4.3 Data Scaling Behaviors of PI-DeepONet and SepONet
We consider the same heat equation benchmark described in Section 3.1 to examine the performance and scalability of SepONet. The training setup for SepONet mirrors that of PI-DeepONet, with the key difference being that SepONet utilizes two trunk nets to independently process the temporal input $t$ and the spatial input $x$. Additionally, SepONet employs a distinct sampling strategy for residual points, as detailed in Section 4.2. We evaluate prediction accuracy on the same 100 test initial conditions, specifically calculating the mean relative $\ell_{2}$ error and RMSE over the test dataset as accuracy metrics.
As shown in Fig. 5, when fixing the number of training functions $N_{f}$ and increasing the number of training points $N_{c}$, or vice versa, the test accuracy of SepONet consistently improves and is superior to that of PI-DeepONet. This is expected since the solution structure in Eq. 16 is separable in terms of different variables, making it easier for SepONet to learn basis functions with different scales for different variables.
Fig. 6 provides a detailed comparison of training time and GPU memory usage for SepONet and PI-DeepONet with respect to data size. As can be seen, when fixing the number of training functions $N_{f}$ and increasing the number of training points $N_{c}$, or vice versa, the training cost for PI-DeepONet significantly increases, whereas the training cost for SepONet remains relatively low. Specifically, for $N_{f}=100$ and $N_{c}=128^{2}$, PI-DeepONet’s training time is approximately 4.15 hours, while SepONet’s is around 0.05 hours. PI-DeepONet’s GPU memory usage exceeds 30 GB, while SepONet’s is under 1 GB. These findings confirm that SepONet scales much better in terms of both training time and memory usage.
Figure 4: A subset of the basis functions learned by SepONet for the heat equation. The first row shows the basis functions $\tau_{k}(t)$ for the temporal variable $t$, and the second row shows the basis functions $\zeta_{k}(x)$ for the spatial variable $x$. These basis functions enable SepONet to efficiently and accurately represent the solution of the heat equation.
Figure 5: Comparison of relative $\ell_{2}$ error and RMSE for PI-DeepONet and SepONet as a function of the number of training points $(N_{c})$ and the number of input training functions $(N_{f})$. When varying one, the other is fixed. The accuracy of SepONet consistently improves with increasing data size and is better than that of PI-DeepONet.
Figure 6: Comparison of training time and GPU memory usage for PI-DeepONet and SepONet as a function of the number of training points $(N_{c})$ and the number of input training functions $(N_{f})$. When varying one, the other is fixed. SepONet shows significantly lower training time and memory usage compared to PI-DeepONet, demonstrating better scalability.
4.4 Universal Approximation Property of SepONet
The universal approximation property of DeepONet has been discussed in [23, 14]. Here we present the universal approximation theorem to show that proposed separable operator networks can also approximate any nonlinear continuous operators that map infinite-dimensional function spaces to others.
Theorem 1 (Universal Approximation Theorem for Separable Operator Networks).
Suppose that $\sigma$ is a Tauber-Wiener function, $g$ is a sinusoidal function, $\mathcal{X}$ is a Banach space, $K\subseteq\mathcal{X}$, $K_{1}\subseteq\boldsymbol{R}^{d_{1}}$ and $K_{2}\subseteq\boldsymbol{R}^{d_{2}}$ are three compact sets in $\mathcal{X}$, $\boldsymbol{R}^{d_{1}}$ and $\boldsymbol{R}^{d_{2}}$, respectively, $\mathcal{U}$ is a compact set in $C\left(K\right)$, $G$ is a nonlinear continuous operator, which maps $\mathcal{U}$ into a compact set $\mathcal{S}\subseteq C\left(K_{1}\times K_{2}\right)$, then for any $\epsilon>0$, there are positive integers $n$, $r$, $m$, constants $c_{i}^{k}$, $\zeta_{k}^{1}$, $\zeta_{k}^{2}$, $\xi_{ij}^{k}$, $\theta_{i}^{k}\in\boldsymbol{R}$, points $\omega_{k}^{1}\in\boldsymbol{R}^{d_{1}}$, $\omega_{k}^{2}\in\boldsymbol{R}^{d_{2}}$, $x_{j}\in K$, $i=1,\ldots,n$, $k=1,\ldots,r$, $j=1,\ldots,m$, such that
$\left|G(u)(y)-\sum_{k=1}^{r}\underbrace{\sum_{i=1}^{n}c_{i}^{k}\sigma\left(\sum_{j=1}^{m}\xi_{ij}^{k}u\left(x_{j}\right)+\theta_{i}^{k}\right)}_{branch}\underbrace{g\left(w_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)}_{trunk_{1}}\underbrace{g\left(w_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)}_{trunk_{2}}\right|<\epsilon$ (22)
holds for all $u\in\mathcal{U}$, $y=(y_{1},y_{2})\in K_{1}\times K_{2}$.
Proof.
The proof can be found in Appendix A.1. ∎
Remark 1.
The definition of the Tauber-Wiener function is given in Appendix A. It is worth noting that many common-used activations, such as ReLU, GELU and Tanh, are Tauber-Wiener functions.
Remark 2.
Here we show the approximation property of a separable operator network with two trunk nets, by repeatedly applying trigonometric angle addition formula, it is trivial to separate $y$ as $(y_{1},y_{2},\ldots,y_{d})\in K_{1}\times K_{2}\times\ldots\times K_{d}$ and extend Theorem (22) to $d$ trunk nets.
Remark 3.
In our assumptions, we restrict the activation function for the trunk nets to be sinusoidal. This choice is motivated by the natural suitability of sinusoidal functions for constructing basis functions [54] and their empirical effectiveness in solving PDEs [55]. However, it would be interesting to explore whether Theorem 1 still holds when $g$ is a more general activation function, such as a Tauber-Wiener function. We will leave this investigation for future work.
5 Numerical Results
In this section, we present a series of numerical studies to demonstrate the expressive power and effectiveness of SepONet on various PDEs, including diffusion-reaction, advection, and Burgers’ equations. We show that SepONet scales significantly better than PI-DeepONet in terms of GPU memory usage (up to 82× memory reduction) and training time (up to 112× faster) using the same training setup.
Across all benchmarks, we apply Tanh activation for the branch net and Sine activation for the trunk nets. Note that all benchmarks are time-dependent 1D PDEs, so there are two trunk nets for SepONet to process $t$ and $x$ separately, while PI-DeepONet uses a single trunk net to process $t$ and $x$ together.
For the sampling strategy, we set $N_{r}=N_{I}^{2}=N_{b}^{2}=N_{c}$ in Eq. (7), where $N_{c}$ is referred to as the number of training points. PI-DeepONet randomly samples initial points $\bigl{\{}(x_{I}^{(j)},0)\bigl{\}}_{j=1}^{N_{I}}$, boundary points $\bigl{\{}(0,t_{b}^{(j)}),(1,t_{b}^{(j)})\bigl{\}}_{j=1}^{N_{b}}$, and residual points $\bigl{\{}(x_{r}^{(j)},t_{r}^{(j)})\bigl{\}}_{j=1}^{N_{r}}$ within the computational domain to enforce the initial, boundary conditions and PDE residual, respectively. SepONet follows the same strategy for sampling initial and boundary points. However, for residual points, SepONet samples $\sqrt{N_{r}}$ points for each axis. Thanks to SepONet’s structure, it can compute evaluations on the entire mesh grid created by these inputs, as detailed in Section 4.2.
Both PI-DeepONet and SepONet are trained by minimizing the physics loss (Eq. 6) using gradient descent with the Adam optimizer [47]. The initial learning rate is $1\times 10^{-3}$ and decays by a factor of 0.9 every 1,000 iterations. Additionally, we resample input training functions and training points (including residual, initial, and boundary points) every 100 iterations. Training for both models is conducted for a fixed number of iterations which varies for different benchmarks. We report the relative $\ell_{2}$ error, root mean squared error (RMSE), GPU memory usage and total training time as metrics to assess the performance of PI-DeepONet and SepONet. Specifically, the mean and standard deviation of the relative $\ell_{2}$ error and RMSE are calculated over all functions in the test dataset.
We note that no extensive hyperparameter tuning was performed for either PI-DeepONet or SepONet. The architectures of the branch and trunk networks, as well as the weight coefficients in the physics loss (Eq. 6), were adopted from the original PI-DeepONet configuration presented in [27]. The code in this study is implemented using JAX and Equinox libraries [49, 50], and all training is performed on a single NVIDIA A100 GPU with 80 GB of memory.
5.1 Diffusion-reaction Systems
Consider the following nonlinear diffusion-reaction system with a source term $u(x)$:
$\frac{\partial s}{\partial t}=D\frac{\partial^{2}s}{\partial x^{2}}+ks^{2}+u(x),\quad(x,t)\in(0,1)\times(0,1]$ (23)
with zero initial and boundary conditions, where $D=0.01$ is the diffusion coefficient and $k=0.01$ is the reaction rate. The goal is to learn the operator $G$ that maps the source term $u(x)$ to the solution $s(x,t)$. We approximate the solution operator using SepONet and PI-DeepONet. The encoder for both models maps the input function $u$ to its point-wise evaluations at 128 equi-spaced sensors in $[0,1]$. The branch net and trunk nets for both models are 5-layer fully-connected neural networks with 50 neurons per hidden layer. We train both models for 50,000 iterations to optimize the physics loss (7) with $\lambda_{I}=\lambda_{b}=1$. The input training source terms are sampled from a mean-zero Gaussian random field (GRF) [56] with a length scale parameter of 0.2. To generate the test dataset, we sample 100 different source terms from the same GRF and apply a second-order implicit finite difference method [57] to obtain the solutions on a uniform $128\times 128$ grid.
Fixing the number of training functions at $N_{f}=100$ and varying the number of training points $N_{c}$, the results are reported in Table 3. As $N_{c}$ increases, the accuracy and robustness of both PI-DeepONet and SepONet improve, as reflected by the decreased mean and standard deviation of the relative $\ell_{2}$ error and RMSE. However, for PI-DeepONet, achieving this $\approx 100$% accuracy improvement results in a significant increase in training cost: the training time increases from 0.0433 hours to 2.8025 hours, and total GPU memory usage rises from 0.729 GB to 35.371 GB. In contrast, SepONet maintains low and stable training costs, demonstrating better scalability and efficiency.
Similarly, Table 4 shows that when $N_{c}=128^{2}$ is fixed and $N_{f}$ varies, the accuracy for both models improves with more training functions. However, PI-DeepONet’s memory usage and training time increase substantially, while SepONet achieves similar or better accuracy with significantly lower training costs. When $N_{f}\leq 10$, the accuracy is low for both models, indicating the necessity of using more training functions for better results.
These results highlight that while both models benefit from increased training data, SepONet scales much better in terms of memory usage and training time. SepONet allows for low-cost training while still achieving good results, whereas PI-DeepONet requires sacrificing training cost for better accuracy and robustness.
Table 3: Performance comparison of PI-DeepONet and SepONet on the diffusion-reaction equation with varying number of training points ($N_{c}$) and fixed number of input training functions ($N_{f}=100$).
Metrics Models \ $N_{c}$ $8^{2}$ $16^{2}$ $32^{2}$ $64^{2}$ $128^{2}$
Relative $\ell_{2}$ error (%) PI-DeepONet 1.39 ± 0.71 1.11 ± 0.59 0.87 ± 0.41 0.83 ± 0.35 0.73 ± 0.34
SepONet 1.49 ± 0.82 0.79 ± 0.35 0.70 ± 0.33 0.62 ± 0.28 0.62 ± 0.26
RMSE (×$\times 10^{-2}$) PI-DeepONet 0.58 ± 0.29 0.46 ± 0.22 0.37 ± 0.20 0.36 ± 0.20 0.32 ± 0.18
SepONet 0.62 ± 0.28 0.35 ± 0.22 0.32 ± 0.23 0.28 ± 0.20 0.29 ± 0.21
Memory (GB) PI-DeepONet 0.729 1.227 3.023 9.175 35.371
SepONet 0.715 0.717 0.715 0.717 0.719
Training time (hours) PI-DeepONet 0.0433 0.0641 0.1497 0.7252 2.8025
SepONet 0.0403 0.0418 0.0430 0.0427 0.0326
Table 4: Performance comparison of PI-DeepONet and SepONet on the diffusion-reaction equation with varying number of input training functions ($N_{f}$) and fixed number of training points ($N_{c}=128^{2}$).
Metrics Models \ $N_{f}$ $5$ $10$ $20$ $50$ $100$
Relative $\ell_{2}$ error (%) PI-DeepONet 34.54 ± 27.83 4.23 ± 2.52 1.72 ± 1.00 0.91 ± 0.46 0.73 ± 0.34
SepONet 22.40 ± 12.30 3.11 ± 1.89 1.19 ± 0.74 0.73 ± 0.32 0.62 ± 0.26
RMSE (×$\times 10^{-2}$) PI-DeepONet 14.50 ± 9.04 1.75 ± 0.90 0.71 ± 0.37 0.40 ± 0.28 0.32 ± 0.18
SepONet 9.34 ± 5.37 1.36 ± 1.07 0.50 ± 0.29 0.34 ± 0.25 0.29 ± 0.21
Memory (GB) PI-DeepONet 2.767 5.105 9.239 17.951 35.371
SepONet 0.719 0.719 0.717 0.717 0.719
Training time (hours) PI-DeepONet 0.1268 0.2218 0.5864 1.4018 2.8025
SepONet 0.0375 0.0390 0.0370 0.0317 0.0326
5.2 Advection Equation
Consider the following linear advection equation parameterized by the variable coefficients $u(x)$:
$\frac{\partial s}{\partial t}+u(x)\frac{\partial s}{\partial x}=0,\quad(x,t)\in(0,1)\times(0,1)$ (24)
with the initial and boundary condition
$s(x,0)=\sin(\pi x),\quad x\in(0,1),$ (25)
$s(0,t)=\sin\left(\frac{\pi}{2}t\right),\quad t\in(0,1).$
The goal is to learn the solution operator $G$ mapping the coefficient $u(x)$ to the corresponding solution $s(x,t)$. We approximate the solution operator using SepONet and PI-DeepONet. The encoder for both models maps the input function $u$ to its point-wise evaluations at 128 equi-spaced sensors in $[0,1]$. The branch net and trunk nets for both models are 6-layer fully-connected neural networks with 100 neurons per hidden layer. We train both models for 120,000 iterations to optimize the physics loss (7) with $\lambda_{I}=\lambda_{b}=100$. The input training functions $u(x)$ are strictly positive by defining $u(x)=v(x)-\min_{x}v(x)+1$, where $v$ is sampled from a GRF with length scale parameter of 0.2. To create the test dataset, we generate 100 new coefficients in the same manner that were not used in training and apply the Lax–Wendroff scheme [57] to solve the advection equation on a uniform $128\times 128$ grid.
Table 5 presents the results for a fixed number of training functions $N_{f}=100$ and varying numbers of training points $N_{c}$. As $N_{c}$ increases, both PI-DeepONet and SepONet show improved accuracy and robustness, consistent with the diffusion-reaction equation findings. However, the advection equation results highlight more pronounced differences in computational demands due to the larger neural networks required. Specifically, PI-DeepONet exhibits a $\approx 105$× increase in training time (from 0.0787 to 8.231 hours) and a $\approx 62$× increase in GPU memory usage (from 0.967 GB to 59.806 GB) to achieve its accuracy improvements. These increases are significantly larger than those observed in the diffusion-reaction equation case, indicating the greater computational challenge for PI-DeepONet with larger neural networks. In contrast, SepONet maintains low and stable training costs across both equation types, emphasizing its computational efficiency and scalability.
Table 6 shows the impact of varying $N_{f}$ while keeping $N_{c}=128^{2}$ fixed. Unlike the diffusion-reaction equation case, where low $N_{f}$ values resulted in poor accuracy for both models, the advection equation results show reasonable accuracy even with fewer training functions. This suggests that the advection equation may be less sensitive to the number of training functions, with SepONet allowing larger $N_{f}$ to improve accuracy without increasing training costs.
These findings reinforce the scalability and efficiency advantages of SepONet and highlight how different partial differential equations and neural network sizes impact model performance and computational requirements. The advection equation results demonstrate that the choice between PI-DeepONet and SepONet is critical for computationally demanding problems, where SepONet’s efficiency leads to substantial resource savings without compromising accuracy.
Table 5: Performance comparison of PI-DeepONet and SepONet on the advection equation with varying number of training points ($N_{c}$) and fixed number of input training functions ($N_{f}=100$).
Metrics Models \ $N_{c}$ $8^{2}$ $16^{2}$ $32^{2}$ $64^{2}$ $128^{2}$
Relative $\ell_{2}$ error (%) PI-DeepONet 9.27 ± 1.94 7.55 ± 1.86 6.79 ± 1.84 6.69 ± 1.95 5.72 ± 1.57
SepONet 14.29 ± 2.65 11.96 ± 2.17 6.14 ± 1.58 5.80 ± 1.57 4.99 ± 1.40
RMSE (×$\times 10^{-2}$) PI-DeepONet 5.88 ± 1.34 4.79 ± 1.27 4.31 ± 1.23 4.24 ± 1.29 3.63 ± 1.05
SepONet 9.06 ± 1.88 7.58 ± 1.55 3.90 ± 1.09 3.69 ± 1.07 3.17 ± 0.95
Memory (GB) PI-DeepONet 0.967 1.741 5.103 17.995 59.806
SepONet 0.713 0.715 0.715 0.715 0.719
Training time (hours) PI-DeepONet 0.0787 0.1411 0.4836 2.3987 8.231
SepONet 0.0843 0.0815 0.0844 0.0726 0.0730
Table 6: Performance comparison of PI-DeepONet and SepONet on the advection equation with varying number of input training functions ($N_{f}$) and fixed number of training points ($N_{c}=128^{2}$).
Metrics Models \ $N_{f}$ $5$ $10$ $20$ $50$ $100$
Relative $\ell_{2}$ error (%) PI-DeepONet 9.64 ± 2.91 8.77 ± 2.23 7.57 ± 1.98 6.69 ± 1.93 5.72 ± 1.57
SepONet 7.62 ± 2.06 6.59 ± 1.71 5.47 ± 1.57 5.18 ± 1.51 4.99 ± 1.40
RMSE (×$\times 10^{-2}$) PI-DeepONet 6.11 ± 1.90 5.55 ± 1.51 4.80 ± 1.33 4.24 ± 1.28 3.63 ± 1.05
SepONet 4.83 ± 1.38 4.18 ± 1.17 3.47 ± 1.06 3.29 ± 1.02 3.17 ± 0.95
Memory (GB) PI-DeepONet 3.021 5.611 9.707 34.511 59.806
SepONet 0.713 0.715 0.719 0.719 0.719
Training time (hours) PI-DeepONet 0.3997 1.0766 1.9765 4.411 8.231
SepONet 0.0754 0.0715 0.0736 0.0720 0.0730
5.3 Burgers’ Equation
Consider the nonlinear Burgers’ equation:
$\frac{\partial s}{\partial t}+s\frac{\partial s}{\partial x}-\nu\frac{\partial^{2}s}{\partial x^{2}}=0,\quad(x,t)\in(0,1)\times(0,1],$ (26)
$s(x,0)=u(x),\quad x\in(0,1)$
with periodic boundary conditions:
$s(0,t)=s(1,t),$ (27)
$\frac{\partial s}{\partial x}(0,t)=\frac{\partial s}{\partial x}(1,t),$
where $t\in(0,1)$ and $\nu=0.01$ is the viscosity. We aim to approximate the solution operator mapping the initial condition $u(x)$ to the solution $s(t,x)$ by PI-DeepONet and SepONet. The encoder for both models maps the initial condition $u$ to its point-wise evaluations at 101 equi-spaced sensors in $[0,1]$. The branch net and trunk nets for both models are 7-layer fully-connected neural networks with 100 neurons per hidden layer. We train both models for 80,000 iterations to optimize the physics loss (7) with $\lambda_{I}=20,\lambda_{b}=1$. Here the input training initial conditions are sampled from a GRF ~ $\mathcal{N}\left(0,25^{2}\left(-\Delta+5^{2}I\right)^{-4}\right)$ using the Chebfun package [48], satisfying the periodic boundary conditions. Synthetic test dataset consists of 100 unseen initial functions and their corresponding solutions, which are generated from the same GRF and are solved by spectral method on a $101\times 101$ uniform grid using the spinOp library [58], respectively.
Tables 8 and 7 present the results for varying numbers of training points $N_{c}$ and training functions $N_{f}$, respectively. These results demonstrate that the Burgers’ equation presents a more significant challenge than the previous examples, necessitating larger $N_{c}$ and $N_{f}$ values to achieve satisfactory accuracy.
PI-DeepONet struggles with this increased complexity, primarily due to memory constraints. As $N_{c}$ increases from $8^{2}$ to $64^{2}$ (with $N_{f}=100$), PI-DeepONet’s relative $\ell_{2}$ error only improves from 29.33% to 13.72%, while its memory usage jumps from 1.253 GB to 18.001 GB. Similarly, when $N_{f}$ increases from 5 to 50 (with $N_{c}=128^{2}$), the error decreases from 28.48% to 12.33%, but memory usage increases from 5.085 GB to 35.433 GB. Critically, PI-DeepONet encounters out-of-memory issues when $N_{c}=128^{2}$ and $N_{f}=100$, preventing further accuracy improvements and highlighting its limitations in handling large-scale problems.
In contrast, SepONet demonstrates superior scalability and efficiency. It not only handles the configurations that cause PI-DeepONet to fail but also allows for further increases in $N_{c}$ and $N_{f}$ to achieve better accuracy. Table 9 illustrates this capability, showing that SepONet can utilize $N_{c}$ up to $512^{2}$ and $N_{f}$ up to 800, achieving a relative $\ell_{2}$ error as low as 4.12%. Remarkably, SepONet maintains relatively low training costs even at these large scales, with a maximum training time of 0.478 hours and memory usage of 10.485 GB for the largest configuration.
These results underscore a critical limitation of PI-DeepONet: its inability to effectively handle large-scale problems that require extensive neural networks to represent the solution operator and demand large $N_{c}$ and $N_{f}$ for improved generalization error. The Burgers’ equation example clearly demonstrates that as the complexity of the PDE increases, PI-DeepONet’s performance suffers due to memory constraints, while SepONet’s architecture allows it to scale efficiently, providing better accuracy with manageable computational resources.
This contrast in scalability and performance highlights SepONet’s significant advantage in tackling complex parametric PDE systems, where the ability to leverage larger datasets and more extensive neural networks is crucial for achieving satisfactory accuracy and generalization.
Table 7: Performance comparison of PI-DeepONet and SepONet on the Burgers’ equation with varying number of training points ($N_{c}$) and fixed number of input training functions ($N_{f}=100$). The ’-’ symbol indicates that results are not available due to out-of-memory issues.
Metrics Models \ $N_{c}$ $8^{2}$ $16^{2}$ $32^{2}$ $64^{2}$ $128^{2}$
Relative $\ell_{2}$ error (%) PI-DeepONet 29.33 ± 3.85 20.31 ± 4.31 14.17 ± 5.25 13.72 ± 5.59 -
SepONet 29.42 ± 3.79 31.53 ± 3.44 28.74 ± 4.11 11.85 ± 4.06 7.51 ± 4.04
RMSE (×$\times 10^{-2}$) PI-DeepONet 4.19 ± 2.79 2.82 ± 1.86 2.23 ± 2.10 2.20 ± 2.13 -
SepONet 4.18 ± 2.74 4.44 ± 2.81 4.11 ± 2.76 1.80 ± 1.60 1.23 ± 1.44
Memory (GB) PI-DeepONet 1.253 2.781 5.087 18.001 -
SepONet 0.603 0.605 0.603 0.605 0.716
Training time (hours) PI-DeepONet 0.1497 0.2375 0.6431 3.2162 -
SepONet 0.0706 0.0719 0.0716 0.0718 0.0605
Table 8: Performance comparison of PI-DeepONet and SepONet on the Burgers’ equation with varying number of input training functions ($N_{f}$) and fixed number of training points ($N_{c}=128^{2}$). The ’-’ symbol indicates that results are not available due to out-of-memory issues.
Metrics Models \ $N_{f}$ $5$ $10$ $20$ $50$ $100$
Relative $\ell_{2}$ error (%) PI-DeepONet 28.48 ± 4.17 28.63 ± 4.10 28.26 ± 4.38 12.33 ± 5.14 -
SepONet 27.79 ± 4.40 28.16 ± 4.24 22.78 ± 6.47 10.25 ± 4.44 7.51 ± 4.04
RMSE (×$\times 10^{-2}$) PI-DeepONet 4.09 ± 2.77 4.11 ± 2.78 4.07 ± 2.78 1.96 ± 1.92 -
SepONet 4.01 ± 2.75 4.05 ± 2.76 3.30 ± 2.55 1.65 ± 1.64 1.23 ± 1.44
Memory (GB) PI-DeepONet 5.085 9.695 17.913 35.433 -
SepONet 0.605 0.607 0.607 0.609 0.716
Training time (hours) PI-DeepONet 0.5135 1.3896 2.6904 5.923 -
SepONet 0.0725 0.0707 0.0703 0.0612 0.0605
Table 9: Additional SepONet results for Burgers’ equation, demonstrating that larger $N_{c}$ and $N_{f}$ can be used to enhance accuracy with minimal cost increase.
Metrics \ $N_{c}$ $N_{f}$ $128^{2}$ 400 $128^{2}$ 800 $256^{2}$ 400 $256^{2}$ 800 $512^{2}$ 400 $512^{2}$ 800
Relative $\ell_{2}$ error (%) 6.60 ± 4.17 6.21± 4.00 5.68 ± 3.54 4.46 ± 3.39 5.38 ± 3.60 4.12 ± 2.67
RMSE (×$\times 10^{-2}$) 1.11 ± 1.40 1.06 ± 1.37 0.95 ± 1.20 0.77 ± 1.77 0.91 ± 1.24 0.69 ± 0.93
Memory (GB) 0.966 1.466 2.466 4.466 5.593 10.485
Training time (hours) 0.0771 0.0957 0.1238 0.1717 0.2751 0.478
6 Discussion
The field of operator learning faces a critical dilemma. Deep Operator Networks (DeepONets) offer fast training but require extensive data generation, which can be prohibitively expensive or impractical for complex systems. Physics-informed DeepONets (PI-DeepONets) relax the data requirement but at the cost of resource-intensive training processes. This creates a challenging trade-off: balancing resource allocation either before training (data generation) or during training (computational resources).
Our proposed Separable Operator Networks (SepONet) effectively address both of these concerns. Inspired by the separation of variables technique typically used for linear PDEs, SepONet constructs its own basis functions to approximate a nonlinear operator. SepONet’s expressive power is guaranteed by the universal approximation property (Theorem 1), ensuring that SepONet can approximate any nonlinear continuous operator with arbitrary accuracy. Our numerical results further support this theoretical guarantee, demonstrating SepONet’s ability to handle complex, nonlinear systems efficiently.
By leveraging independent trunk nets for different variables, SepONet allows for an efficient implementation via forward-mode automatic differentiation (AD). This approach achieves remarkable efficiency in both data utilization and computational resources. Specifically, SepONet does not require ground truth PDE solutions as it is trained solely via optimizing the physics loss, leading to significant data efficiency by eliminating the need for expensive simulations. In terms of computational resources, SepONet maintains stable GPU memory usage and training time, even when increasing training data and network size, in contrast to PI-DeepONet’s dramatic resource consumption increases under similar scaling.
However, SepONet has certain limitations that warrant further research and development. The mesh grid structure of SepONet’s solution, while enabling efficient training through forward-mode AD, may limit its flexibility in handling PDEs with irregular geometries. Addressing this limitation could involve developing adaptations or hybrid approaches that accommodate more complex spatial domains [59, 60, 61], potentially expanding SepONet’s applicability to a broader range of physical problems. Additionally, while the linear decoder allows for an efficient SepONet implementation, a very large number of basis functions may be needed for accurate linear representation in some problems [39]. Developing a nonlinear decoder version of SepONet will be useful to balance accuracy and efficient training. Moreover, implementing an adaptive weighting strategy [62, 63, 64] for different loss terms in the physics loss function, instead of using predefined fixed weights, could lead to improved accuracy and faster convergence. Finally, empirical observations suggest that training accuracy and robustness improve with an increase in input training functions. However, the neural scaling laws in physics-informed operator learning remain unexplored, presenting an intriguing theoretical challenge for future investigation.
Acknowledgments
We would like to thank Wolfger Peelaers for the valuable discussions and insightful comments.
Appendix A Universal Approximation Theorem for Separable Operator Networks
Here we present the universal approximation theorem for the proposed separable operator networks 1, review well-known theoretical results of approximating continuous functions and functionals, and show the preliminary lemmas and proofs for Theorem 1. We refer our readers to [23, 65] for detailed proofs of Theorems 2, 3, 4. Main notations are listed in Table 10.
Definition 1 (TW).
If a function $g:\boldsymbol{R}\rightarrow\boldsymbol{R}$ (continuous or discontinuous) satisfies that all the linear combinations $\sum_{i=1}^{N}c_{i}g\left(\lambda_{i}x+\theta_{i}\right)$, $\lambda_{i}\in\boldsymbol{R}$, $\theta_{i}\in\boldsymbol{R}$, $c_{i}\in\boldsymbol{R}$, $i=1,2,\ldots,N$, are dense in every $C[a,b]$, then $g$ is called a Tauber-Wiener (TW) function.
Remark 1 (Density in $C[a,b]$).
A set of functions is said to be dense in $C[a,b]$ if every function in the space of continuous functions on the interval $[a,b]$ can be approximated arbitrarily closely by functions from the set.
Definition 2 (Compact Set).
Suppose that $X$ is a Banach space, $V\subseteq X$ is called a compact set in $X$, if for every sequence $\left\{x_{n}\right\}_{n=1}^{\infty}$ with all $x_{n}\in V$, there is a subsequence $\left\{x_{n_{k}}\right\}$, which converges to some element $x\in V$.
Theorem 1 (Universal Approximation Theorem for Separable Operator Networks).
Suppose that $\sigma\in(\mathrm{TW})$, $g$ is a sinusoidal function, $X$ is a Banach Space, $K\subseteq X$, $K_{1}\subseteq\boldsymbol{R}^{d_{1}}$ and $K_{2}\subseteq\boldsymbol{R}^{d_{2}}$ are three compact sets in $X$, $\boldsymbol{R}^{d_{1}}$ and $\boldsymbol{R}^{d_{2}}$, respectively, $\mathcal{U}$ is a compact set in $C\left(K\right)$, $G$ is a nonlinear continuous operator, which maps $\mathcal{U}$ into a compact set $\mathcal{S}\subseteq C\left(K_{1}\times K_{2}\right)$, then for any $\epsilon>0$, there are positive integers $n$, $r$, $m$, constants $c_{i}^{k}$, $\zeta_{k}^{1}$, $\zeta_{k}^{2}$, $\xi_{ij}^{k}$, $\theta_{i}^{k}\in\boldsymbol{R}$, points $\omega_{k}^{1}\in\boldsymbol{R}^{d_{1}}$, $\omega_{k}^{2}\in\boldsymbol{R}^{d_{2}}$, $x_{j}\in K_{1}$, $i=1,\ldots,n$, $k=1,\ldots,r$, $j=1,\ldots,m$, such that
$\left|G(u)(y)-\sum_{k=1}^{r}\sum_{i=1}^{n}c_{i}^{k}\sigma\left(\sum_{j=1}^{m}\xi_{ij}^{k}u\left(x_{j}\right)+\theta_{i}^{k}\right)g\left(w_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)g\left(w_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)\right|<\epsilon$ (28)
holds for all $u\in\mathcal{U}$, $y=(y_{1},y_{2})\in K_{1}\times K_{2}$.
Theorem 2 ([23]).
Suppose that $K$ is a compact set in $\boldsymbol{R}^{n}$, $\mathcal{S}$ is a compact set in $C(K)$, $g\in(\mathrm{TW})$, then for any $\epsilon>0$, there exist a positive integer $N$, real numbers $\theta_{i}$, vectors $\omega_{i}\in\boldsymbol{R}^{n}$, $i=1,\ldots,N$, which are independent of $f\in C(K)$ and constants $c_{i}(f)$, $i=1,\ldots,N$ depending on $f$, such that
$\left|f(x)-\sum_{i=1}^{N}c_{i}(f)g\left(\omega_{i}\cdot x+\theta_{i}\right)\right|<\epsilon$ (29)
holds for all $x\in K$ and $f\in\mathcal{S}$. Moreover, each $c_{i}(f)$ is a linear continuous functional defined on $\mathcal{S}$.
Theorem 3 ([23]).
Suppose that $\sigma\in(\mathrm{TW}),X$ is a Banach Space, $K\subseteq X$ is a compact set, $\mathcal{U}$ is a compact set in $C(K)$, $f$ is a continuous functional defined on $\mathcal{U}$, then for any $\epsilon>0$, there are positive integers $N$, $m$ points $x_{1},\ldots,x_{m}\in K$, and real constants $c_{i}$, $\theta_{i}$, $\xi_{ij}$, $i=1,\ldots,N$, $j=1,\ldots,m$, such that
$\left|f(u)-\sum_{i=1}^{N}c_{i}\sigma\left(\sum_{j=1}^{m}\xi_{ij}u\left(x_{j}\right)+\theta_{i}\right)\right|<\epsilon$ (30)
holds for all $u\in\mathcal{U}$.
Theorem 4 (Weierstrass Approximation Theorem [65]).
Suppose $f\in C[a,b],$ then for every $\epsilon>0,$ there exists a polynomial p such that for all $x$ in $[a,b]$, we have $|f(x)-p(x)|<\epsilon.$
Table 10: Notations and Symbols
$X$ some Banach space with norm $\|\cdot\|_{X}$
$\boldsymbol{R}^{d}$ Euclidean space of dimension $d$
$K$ some compact set in a Banach space
$C(K)$ Banach space of all continuous functions defined on $K$, with norm $\|f\|_{C(K)}=\max_{x\in K}|f(x)|$
$\tilde{C}[a,b]$ the space of functions in $C[a,b]$ satisfying $f(a)=f(b)$
$V$ some compact set in $C(K)$
$u(x)$ some input function
$\mathcal{U}$ the space of input functions
$G$ some continuous operator
$G(u)(y)$ or $s(y)$ some output function that is mapped from the corresponding input function $u$ by the operator $G$
$\mathcal{S}$ the space of output functions
(TW) all the Tauber-Wiener functions
$\sigma$ and $g$ activation function for branch net and trunk nets in Theorems 1
$\left\{x_{1},x_{2},\ldots,x_{m}\right\}$ $m$ sensor points for identifying input function $u$
$r$ rank of some deep operator network or separable operator network
$n,m$ operator network size hyperparameters in Theorems 1
Corollary 1.1.
Trigonometric polynomials are dense in the space of continuous and periodic functions $\tilde{C}[0,2\pi]:=\{f\in C[0,2\pi]~{}|~{}f(0)=f(2\pi)\}$.
Proof.
For any $\tilde{f}\in\tilde{C}[0,2\pi]$, extend it to a $2\pi$-periodic and continuous function $f$ defined on $\boldsymbol{R}$. It suffices to show that there exists a trigonometric polynomial that approximates $f$ within any $\epsilon>0$. We construct the continuous even functions of $2\pi$ period $g$ and $h$ as:
$g(\theta)=\frac{f(\theta)+f(-\theta)}{2}\quad\text{and}\quad h(\theta)=\frac{f(\theta)-f(-\theta)}{2}\sin(\theta).$ (31)
Let $\phi(x)=g(\arccos x)$ and $\psi(x)=h(\arccos x)$. Since $\phi,\psi$ are continuous functions on $[-1,1]$, by the Weierstrass Approximation Theorem 4, for any $\epsilon>0$, there exist polynomials $p$ and $q$ such that
$|\phi(x)-p(x)|<\frac{\epsilon}{4}\quad\text{and}\quad|\psi(x)-q(x)|<\frac{\epsilon}{4}$ (32)
holds for all $x\in[-1,1]$. Let $x=\cos\theta$, it follows that
$|g(\theta)-p(\cos\theta)|<\frac{\epsilon}{4}\quad\text{and}\quad|h(\theta)-q(\cos\theta)|<\frac{\epsilon}{4}$ (33)
for $\theta\in[0,\pi]$. Because $g$, $h$ and $cosine$ are even and $2\pi$-periodic, (33) holds for all $\theta\in\boldsymbol{R}$. From the definitions of $g$ and $h$, and the fact $\left|\sin\theta\right|\leq 1$, $\left|\sin^{2}\theta\right|\leq 1$, we have
$\left|\frac{f(\theta)+f(-\theta)}{2}\sin^{2}\theta-p(\cos\theta)\sin^{2}\theta\right|<\frac{\epsilon}{4}\quad\text{and}\quad\left|\frac{f(\theta)-f(-\theta)}{2}\sin^{2}\theta-q(\cos\theta)\sin\theta\right|<\frac{\epsilon}{4}.$ (34)
Using the triangle inequality, we obtain
$\left|f(\theta)\sin^{2}\theta-\left[p(\cos\theta)\sin^{2}\theta+q(\cos\theta)\sin\theta\right]\right|<\frac{\epsilon}{2}.$ (35)
Applying the same analysis to
$\tilde{g}(\theta)=\frac{f(\theta+\frac{\pi}{2})+f(-\theta+\frac{\pi}{2})}{2}\quad\text{and}\quad\tilde{h}(\theta)=\frac{f(\theta+\frac{\pi}{2})-f(-\theta+\frac{\pi}{2})}{2}\sin(\theta),$ (36)
we can find polynomials $r$ and $s$ such that
$\left|f\left(\theta+\frac{\pi}{2}\right)\sin^{2}\theta-\left[r(\cos\theta)\sin^{2}\theta+s(\cos\theta)\sin\theta\right]\right|<\frac{\epsilon}{2}$ (37)
holds for all $\theta$. Substituting $\theta$ with $\theta-\frac{\pi}{2}$ gives
$\left|f(\theta)\cos^{2}\theta-\left[r(\sin\theta)\cos^{2}\theta-s(\sin\theta)\cos\theta\right]\right|<\frac{\epsilon}{2}.$ (38)
By the triangle inequality, combining (38) and (35) gives
$\left|f(\theta)-\left[r(\sin\theta)\cos^{2}\theta-s(\sin\theta)\cos\theta+p(\cos\theta)\sin^{2}\theta+q(\cos\theta)\sin\theta\right]\right|<\epsilon$ (39)
holds for all $\theta$. Thus, the trigonometric polynomial
$r(\sin\theta)\cos^{2}\theta-s(\sin\theta)\cos\theta+p(\cos\theta)\sin^{2}\theta+q(\cos\theta)\sin\theta$ (40)
is an $\epsilon$-approximation to $f$. ∎
Remark 2.
If $p(x)$ is a polynomial, it is easy to verify that $p(\cos\theta)$ is a trigonometric polynomial due to the fact $\cos^{n}\theta=\sum_{k=0}^{n}\frac{\binom{n}{k}}{2^{n}}\cos\left((n-2k)\theta\right)$.
Prior to proving Theorem 1, we need to establish the following lemmas.
Lemma 1.
Sine is a Tauber-Wiener function.
Proof.
Assuming the interval to be $[0,\pi]$ first. For every continuous function $f$ on $[0,\pi]$ and any $\epsilon>0$, we can extend $f$ to a continuous function $F$ on $[0,2\pi]$ so that $F(x)=f(x)$ on $[0,\pi]$ and $F(2\pi)=F(0)$. By Corollary 1.1, there exists a trigonometric polynomial
$p(x)=a_{0}+\sum_{n=1}^{N}a_{n}\cos(nx)+b_{n}\sin(nx)$ (41)
such that
$\sup_{x\in[0,\pi]}|f(x)-p(x)|\leq\sup_{x\in[0,2\pi]}|F(x)-p(x)|<\epsilon.$ (42)
Let $c_{0}=a_{0}$, $\lambda_{0}=0$, $\theta_{0}=\frac{\pi}{2}$, $c_{2n-1}=b_{n}$, $\lambda_{2n-1}=n$, $\theta_{2n-1}=0$, $c_{2n}=a_{n}$, $\lambda_{2n}=n$, $\theta_{2n}=\frac{\pi}{2}$, for $n=1,2,\ldots,N$, $p(x)$ is redefined as
$p(x)=\sum_{i=0}^{2N}c_{i}\sin\left(\lambda_{i}x+\theta_{i}\right).$ (43)
Thus we have
$\left|f(x)-\sum_{i=0}^{2N}c_{i}\sin\left(\lambda_{i}x+\theta_{i}\right)\right|<\epsilon$ (44)
for $x\in[0,\pi]$. Now consider a continuous function $g$ on $[a,b]$, define $f(x)\in C[0,\pi]:=g\left(\frac{b-a}{\pi}x+a\right)$, then by (44), we have
$\left|g(x)-\sum_{i=0}^{2N}c_{i}\sin\left(\frac{\pi\lambda_{i}}{b-a}x-\frac{\pi\lambda_{i}a}{b-a}+\theta_{i}\right)\right|<\epsilon$ (45)
holds for all $x\in[a,b]$. Therefore, it follows that for any continuous function $g$ on $[a,b]$ and any $\epsilon>0$, we can approximate $g$ within $\epsilon$ by choosing $N$ sufficiently large and adjusting $c_{i}$, $\lambda_{i}$, $\theta_{i}$ accordingly. Hence, the set of all such linear combinations of $\sin(x)$ is dense in $C[a,b]$, confirming that $\sin(x)$ is a Tauber-Wiener function. ∎
Remark 3.
It is straightforward to conclude that all sinusoidal functions are Tauber-Wiener functions.
Lemma 2.
Suppose that $V_{1}\subseteq X_{1}$, $V_{2}\subseteq X_{2}$ are two compact sets in Banach spaces $X_{1}$ and $X_{2}$, respectively, then their Cartesian product $V_{1}\times V_{2}$ is also compact.
Proof.
For every sequence $\left\{x_{n}^{1},x_{n}^{2}\right\}$ in $V_{1}\times V_{2}$, since $V_{1}$ is compact, $\left\{x_{n}^{1}\right\}$ has a subsequence $\left\{x_{n_{k}}^{1}\right\}$ that converges to some element $x^{1}\in V_{1}$. As well, since $V_{2}$ is compact, there exists a subsequence $\left\{x_{n_{k}}^{2}\right\}$ that converges to $x^{2}\in V_{2}$. It follows that $\left\{x_{n}^{1},x_{n}^{2}\right\}$ converges to $\left(x^{1},x^{2}\right)\in V_{1}\times V_{2}$, thus $V_{1}\times V_{2}$ is compact. ∎
Lemma 3.
Suppose that $X$ is a Banach space, $K_{1}\subseteq X_{1}$, $K_{2}\subseteq X_{2}$ are two compact sets in $X_{1}$ and $X_{2}$, respectively. $\mathcal{U}$ is a compact set in $C(K_{1})$, then the range $G(\mathcal{U})$ of the continuous operator $G$ from $\mathcal{U}$ to $C(K_{2})$ is compact in $C(K_{2})$.
Proof.
For every sequence $\{f_{n}\}$ in $\mathcal{U}$, since $\mathcal{U}$ is compact, there exists a subsequence $\{f_{n_{k}}\}$ that converges to some function $f\in\mathcal{U}$. Since $G$ is continuous, the convergence $f_{n_{k}}\to f$ in $C(K_{1})$ implies
$G(f_{n_{k}})\to G(f)\quad\text{in}\quad C(K_{2}).$ (46)
Thus, for every sequence $\{G(f_{n})\}$ in $G(\mathcal{U})$, there exists a subsequence $\{G(f_{n_{k}})\}$ that converges to $G(f)\in G(\mathcal{U})$. Therefore, the range $G(\mathcal{U})$ of the continuous operator $G$ is compact in $C(K_{2})$. ∎
A.1 Proof of Theorem 1
Proof.
Without loss of generality, we can assume that $g$ is sine function, by Lemma 1, we have $g\in(TW)$; From the assumption that $K_{1}$ and $K_{2}$ are compact, by Lemma 2, $K_{1}\times K_{2}$ is compact; Since $G$ is a continuous operator that maps $\mathcal{U}$ into $C(K_{1}\times K_{2})$, it follows that the range $G(\mathcal{U})=\left\{G(u):u\in\mathcal{U}\right\}$ is compact in $C(K_{1}\times K_{2})$ due to Lemma 3; Thus by Theorem 2, for any $\epsilon>0$, there exists a positive integer $N$, real numbers $c_{k}(G(u))$ and $\zeta_{k}$, vectors $\omega_{k}\in R^{d_{1}+d_{2}}$, $k=1,\ldots,N$, such that
$\left|G(u)(y)-\sum_{k=1}^{N}c_{k}(G(u))g\left(\omega_{k}\cdot y+\zeta_{k}\right)\right|<\frac{\epsilon}{2}$ (47)
holds for all $y\in K_{1}\times K_{2}$ and $u\in C(K)$. Let $(\omega_{k}^{1},\omega_{k}^{2})=\omega_{k}$, where $\omega_{k}^{1}\in\boldsymbol{R}^{d_{1}}$ and $\omega_{k}^{2}\in\boldsymbol{R}^{d_{2}}$. Utilizing the trigonometric angle addition formula, we have
$g\left(\omega_{k}\cdot y+\zeta_{k}\right)=g\left(\omega_{k}^{1}\cdot y_{1}+\zeta_{k}\right)g\left(\omega_{k}^{2}\cdot y_{2}+\frac{\pi}{2}\right)+g\left(\omega_{k}^{1}\cdot y_{1}+\zeta_{k}+\frac{\pi}{2}\right)g\left(\omega_{k}^{2}\cdot y_{2}\right).$ (48)
Let $r=2N$, $c_{N+k}(G(u))=c_{k}(G(u))$, $\omega_{N+k}^{1}=\omega_{k}^{1}$, $\omega_{N+k}^{2}=\omega_{k}^{2}$, $\zeta_{k}^{1}=\zeta_{k}$, $\zeta_{k}^{2}=\frac{\pi}{2}$ for $k=1,\ldots,N$, and let $\zeta_{k}^{1}=\zeta_{k}+\frac{\pi}{2}$, $\zeta_{k}^{2}=0$ for $k=N+1,\ldots,r$, (47) can be expressed as:
$\left|G(u)(y)-\sum_{k=1}^{r}c_{k}(G(u))g\left(\omega_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)g\left(\omega_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)\right|<\frac{\epsilon}{2}.$ (49)
Since $G$ is a continuous operator, according to the last proposition of Theorem 2, we conclude that for each $k=1,\ldots,2N$, $c_{k}(G(u))$ is a continuous functional defined on $\mathcal{U}$. Repeatedly applying Theorem 3, for each $k=1,\ldots,2N$, $c_{k}(G(u))$, we can find positive integers $n_{k}$,$m_{k}$, constants $c_{i}^{k}$, $\xi_{ij}^{k}$, $\theta_{i}^{k}\in R$ and $x_{j}\in K_{1}$, $i=1,\ldots,n_{k}$, $j=1,\ldots,m_{k}$, such that
$\left|c_{k}(G(u))-\sum_{i=1}^{n_{k}}c_{i}^{k}\sigma\left(\sum_{j=1}^{m_{k}}\xi_{ij}^{k}u(x_{j})+\theta_{i}^{k}\right)\right|<\frac{\epsilon}{2L}$ (50)
holds for all $k=1,\ldots,r$ and $u\in\mathcal{U}$, where
$L=\sum_{k=1}^{r}\sup_{y_{1}\in K_{2},y_{2}\in K_{3}}\left|g\left(\omega_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)g\left(\omega_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)\right|.$ (51)
Substituting (50) into (49), we obtain that
$\left|G(u)(y)-\sum_{k=1}^{r}\sum_{i=1}^{n_{k}}c_{i}^{k}\sigma\left(\sum_{j=1}^{m_{k}}\xi_{ij}^{k}u\left(x_{j}\right)+\theta_{i}^{k}\right)g\left(w_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)g\left(w_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)\right|<\epsilon$ (52)
holds for all $u\in\mathcal{U}$, $y_{1}\in K_{1}$ and $y_{2}\in K_{2}$. Let $n=\max_{k}{n_{k}}$, $m=\max_{k}{m_{k}}$. For all $n_{k}<i\leq n$, let $c_{i}^{k}=0$. For all $m_{k}<j\leq m$, let $\xi_{ij}^{k}=0$. Then (52) can be rewritten as:
$\left|G(u)(y)-\sum_{k=1}^{r}\sum_{i=1}^{n}c_{i}^{k}\sigma\left(\sum_{j=1}^{m}\xi_{ij}^{k}u\left(x_{j}\right)+\theta_{i}^{k}\right)g\left(w_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)g\left(w_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)\right|<\epsilon,$ (53)
which holds for all $u\in\mathcal{U}$, $y_{1}\in K_{1}$ and $y_{2}\in K_{2}$. This completes the proof of Theorem 1.
Appendix B Basis Functions in SepONet
In this section, we compare the basis functions learned by SepONet with those obtained through the separation of variables method. Recall that the truncated analytical solution at rank $r$ for the heat equation (8) is given by:
$s(x,t)=\sum_{k=1}^{r}A_{k}e^{-k^{2}t}\sin(k\pi x)$ (54)
where the coefficients $A_{k}$ are determined by the initial condition $u(x)$:
[…]
In SepONet, the solution operator $G$ that maps the initial condition to the corresponding solution is approximated by:
$G_{\boldsymbol{\theta}}(u)(x,t)=\sum_{k=1}^{r}b_{k}(u(x_{1}),u(x_{2}),\ldots,u(x_{128}))\tau_{k}(t)\zeta_{k}(x),$ (56)
where $x_{1},x_{2},\ldots,x_{128}$ are 128 equi-spaced sensors in $[0,1]$, $b_{k}$ is the $k$-th output of the branch net, and the basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ are the $k$-th outputs of two separate trunk nets.
Impact of the Rank $r$.
Since $\tau_{k}(t)$ and $\zeta_{k}(x)$ are independent of the initial condition $u$, learning an expressive and rich set of basis functions is crucial for SepONet to generalize to unseen initial conditions. To investigate the impact of the rank $r$ on the generalization error, we used the same training setup and test input functions as in Section 4.3 and trained SepONet with varying rank $r$ from 1 to 50. The mean RMSE between SepONet predictions and the reference solutions over 100 unseen test initial conditions is reported. For comparison, we also calculate the mean RMSE of the truncated analytical solution at rank $r$ for $r$ from 1 to 15. The results are shown in Fig. 7.
As $r$ increases, the truncated analytical solution quickly converges to the reference solution. The nonzero RMSE is caused by numerical errors when computing $A_{k}$ and the fact that the reference solution is not exact due to numerical errors of the forward Euler method. For SepONet, we observed that when $r=1,2$, the mean RMSE matches the truncated solution. However, beyond that point, the error decays more slowly and converges around $r=50$. This indicates that SepONet does not necessarily learn the exact same basis functions as those in Eq. (54). Instead, a larger $r$ is required for SepONet to construct its own set of basis functions, achieving similar accuracy to the truncated solution. We visualize the learned basis functions set of SepONet for different $r$ in Figures 8 to 12.
As can be seen, when $r=1$, SepONet learns the same basis functions as the first term of the truncated solution. When $r=2$, SepONet learns very similar basis functions as the first two terms of the truncated solution. However, when $r=5$, the learned basis functions are not the same as the truncated solution series, but some spatial basis functions still resemble sinusoidal functions and all temporal basis functions are monotonic. As $r$ increases further, the accuracy of SepONet continues to improve, although the learned basis functions increasingly differ from the truncated solution series. This confirms our expectation that SepONet can accurately approximate the solution using its own learned basis functions.
Figure 7: Comparison of RMSE between the truncated analytical solution and SepONet predictions for varying rank $r$. The truncated analytical solution quickly converges, while SepONet shows a slower decay in error, converging around $r=50$.
Figure 8: Learned basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ for $r=1$. SepONet learns the same basis functions as the first term of the truncated solution.
Figure 9: Learned basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ for $r=2$. SepONet learns very similar basis functions as the first two terms of the truncated solution.
Figure 10: Learned basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ for $r=5$.
Figure 11: Learned basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ for $r=10$.
Figure 12: Learned basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ for $r=50$.