### v2

Abstract
Operator learning has become a powerful tool in machine learning for modeling complex physical systems governed by partial differential equations (PDEs). Although Deep Operator Networks (DeepONet) show promise, they require extensive data acquisition. Physics-informed DeepONets (PI-DeepONet) mitigate data scarcity but suffer from inefficient training processes. We introduce Separable Operator Networks (SepONet), a novel framework that significantly enhances the efficiency of physics-informed operator learning. SepONet uses independent trunk networks to learn basis functions separately for different coordinate axes, enabling faster and more memory-efficient training via forward-mode automatic differentiation. We provide a universal approximation theorem for SepONet proving that it generalizes to arbitrary operator learning problems, and then validate its performance through comprehensive benchmarking against PI-DeepONet. Our results demonstrate SepONet’s superior performance across various nonlinear and inseparable PDEs, with SepONet’s advantages increasing with problem complexity, dimension, and scale. For 1D time-dependent PDEs, SepONet achieves up to 112× faster training and 82× reduction in GPU memory usage compared to PI-DeepONet, while maintaining comparable accuracy. For the 2D time-dependent nonlinear diffusion equation, SepONet efficiently handles the complexity, achieving a 6.44% mean relative $\ell_{2}$ test error, while PI-DeepONet fails due to memory constraints. This work paves the way for extreme-scale learning of continuous mappings between infinite-dimensional function spaces. Open source code is available at https://github.com/HewlettPackard/separable-operator-networks.
keywords:
Partial differential equations, separation of variables, scientific machine learning, operator learning, deep operator network (DeepONet), physics-informed neural network
{MSCcodes}
35B30, 35C09, 25G61, 47N20
1 Introduction
Figure 1: Separable operator network (SepONet) architecture for 2D problem instance. A coordinate grid of collocation points $(x^{(i)},y^{(j)})$ can be evaluated efficiently by separating the coordinate axes, feeding them through independent trunk networks, and combining the outputs by outer product to obtain multiple basis function maps. Meanwhile, the branch network processes input functions and outputs coefficients, which are then used to scale and combine the trunk network basis functions by product and sum. Spatiotemporal derivatives of the output predictions are obtained efficiently by forward automatic differentiation due to the independence of trunk networks along each coordinate axis.
Operator learning, which aims to learn mappings between infinite-dimensional function spaces, has gained significant attention in scientific machine learning thanks to its ability to model complex dynamics in physics systems. This approach has been successfully applied to a wide range of applications, including climate modeling [18, 42], multiphysics simulation [32, 4, 37, 30, 23], inverse design [36, 11] and more [17, 15, 49, 12]. Various operator learning algorithms [34, 28, 27, 60, 40, 45, 29, 58, 1] have been developed to address these applications, with Deep Operator Networks (DeepONets) [34] being particularly notable due to their universal approximation guarantee for operators [5] and robustness [35].
To approximate the function operator $G:\mathcal{U}\rightarrow\mathcal{S}$, DeepONets are usually trained in a supervised manner using a dataset of $N_{f}$ pair functions $\left\{u^{(i)},s^{(i)}\right\}_{i=1}^{N_{f}}$, where in the context of parametric partial differential equations (PDEs), each $u^{(i)}$ represents a PDE configuration function and each $s^{(i)}$ represents the corresponding solution. Unlike traditional numerical methods which require repeated simulations for each different PDE configuration, once well trained, a DeepONet allows for efficient parallel inference in abstract infinite-dimensional function spaces. Given new PDE configurations, it can immediately provide the corresponding solutions. However, this advantage comes with a significant challenge: to achieve satisfactory generalization error, the number of required input-output function pairs grows exponentially [34, 24, 31]. Generating enough function pairs can be computationally expensive or even impractical in some applications, creating a bottleneck in the effective deployment of DeepONets.
Physics-informed deep operator networks (PI-DeepONet) [53] have been introduced as a solution to the costly data acquisition problem. Inspired by physics-informed neural networks (PINNs) [44], PI-DeepONet learns operators by constraining the DeepONet outputs to approximately satisfy the underlying governing PDE system parameterized by the input function $u$. This is achieved by penalizing the physics loss (including PDE residual loss, initial loss and boundary loss), thus eliminating the need for ground-truth output functions $s$. However, PI-DeepONet shares the same disadvantage as PINNs in that the training process is both memory-intensive and time-consuming [13, 7]. This inefficiency, common to both PI-DeepONet and PINNs, arises from the need to compute high-order derivatives of the PDE solution with respect to numerous collocation points during physics loss optimization. This computation typically relies on reverse-mode automatic differentiation [2], involving the backpropagation of physics loss through the unrolled computational graph to update model parameters. The inefficiency is even more pronounced for PI-DeepONet, as it requires evaluating the physics loss across multiple PDE configurations. While numerous studies [6, 56, 13, 59, 14, 33, 7, 14] have proposed methods to enhance PINN training efficiency, there has been limited research focused on improving the training efficiency of PI-DeepONet specifically.
To address the inefficiency in training PI-DeepONet, we propose Separable Operator Networks (SepONet), inspired by the separation of variables technique in solving PDEs and recent work on separable PINN [7]. SepONet employs $d$ independent trunk nets to learn basis functions separately for different axes, allowing for predictions of output functions at $N^{d}$ collocation points using only $dN$ inputs, where $d$ is the dimension and $N$ is the number of points sampled along each coordinate axis. This is in contrast to PI-DeepONet which uses only a single trunk net to evaluate predictions along each of $N^{d}$ collocation points individually. The SepONet architecture for a $d=2$ problem instance is shown in Fig. 1. This simple yet effective modification enables fast training and memory-efficient implementation of SepONet by leveraging forward-mode automatic differentiation [19]. Our key contributions are:
  1. 1.
    We introduce SepONet, a physics-informed operator learning framework that significantly enhances scaling efficiency in terms of training time and GPU memory usage, enabling extreme-scale learning of continuous mappings between infinite-dimensional function spaces.
  2. 2.
    We provide a theoretical foundation for SepONet through the universal approximation theorem, demonstrating its capability to approximate any nonlinear continuous operator with arbitrary accuracy.
  3. 3.
    We validate our theoretical results through benchmarking SepONet against PI-DeepONet on a range of 1D and 2D time-dependent PDEs. Our findings reveal that scaling up SepONet in both number of functions and collocation points consistently improves its accuracy, while typically outperforming PI-DeepONet. On 1D time-dependent PDEs, we achieve up to 112× training speed-up with minimal memory increase. Notably, at moderately large scales where PI-DeepONet exhausts 80GB of GPU memory, SepONet operates efficiently with less than 1GB. Furthermore, we observe efficient scaling of SepONet with problem dimension, enabling accurate prediction of 2D time-dependent PDEs at scales where PI-DeepONet fails. The code used to generate the results presented in this manuscript is available at https://github.com/HewlettPackard/separable-operator-networks.
2 Preliminaries
2.1 Operator Learning for Solving Parametric Partial Differential Equations
Let $\mathcal{X}$ and $\mathcal{Y}$ be Banach spaces, with $K\subseteq\mathcal{X}$ and $K_{1}\subseteq\mathcal{Y}$ being compact sets. Consider a nonlinear continuous operator $G:\mathcal{U}\rightarrow\mathcal{S}$, mapping functions from one infinite-dimensional space to another, where $\mathcal{U}\subseteq C(K)$ and $\mathcal{S}\subseteq C(K_{1})$. The goal of operator learning is to approximate the operator $G$ using a model parameterized by $\boldsymbol{\theta}$, denoted as $G_{\boldsymbol{\theta}}$. Here, $\mathcal{U}$ and $\mathcal{S}$ represent spaces of functions where the input and output functions have dimensions $d_{u}$ and $d_{s}$, respectively. We focus on the scalar case where $d_{u}=d_{s}=1$ throughout this paper without loss of generality, however, it should be noted that the results apply to arbitrary $d_{u}$ and $d_{s}$.
In the context of solving parametric partial differential equations (PDEs), consider PDEs of the form:
(1) $\mathcal{N}(u,s)$ $=0,\;\mathcal{I}(u,s)=0,\;\mathcal{B}(u,s)=0,$
where $\mathcal{N}$ is a nonlinear differential operator, $\mathcal{I}$ and $\mathcal{B}$ represent the initial and boundary conditions, $u\in\mathcal{U}$ denotes the PDE configurations (source terms, coefficients, initial conditions, and etc.), and $s\in\mathcal{S}$ denotes the corresponding PDE solution. Assuming that for any $u\in\mathcal{U}$ there exists a unique solution $s\in\mathcal{S}$, we can define the solution operator $G:\mathcal{U}\rightarrow\mathcal{S}$ as $s=G(u)$.
A widely used framework for approximating such an operator $G$ involves constructing $G_{\boldsymbol{\theta}}$ through three maps [25]:
(2) $G_{\boldsymbol{\theta}}\approx G:=\mathcal{D}\circ\mathcal{A}\circ\mathcal{E}.$
First, the encoder $\mathcal{E}:\mathcal{U}\rightarrow\mathbb{R}^{m}$ maps an input function $u\in\mathcal{U}$ to a finite-dimensional feature representation. Next, the approximator $\mathcal{A}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{r}$ transforms this encoded data within the finite-dimensional space $\mathbb{R}^{m}$ to another finite-dimensional space $\mathbb{R}^{r}$. Finally, the decoder $\mathcal{D}:\mathbb{R}^{r}\rightarrow\mathcal{S}$ produces the output function $s(y)=G(u)(y)$ for $y\in K_{1}$.
2.2 Deep Operator Networks (DeepONet)
The original DeepONet formulation [34] can be analyzed through the 3-step approximation framework Eq. 2. The encoder $\mathcal{E}:\mathcal{U}\rightarrow\mathbb{R}^{m}$ maps the input function $u$ to its point-wise evaluations at $m$ fixed sensors $x_{1},x_{2},\ldots,x_{m}\in K$, e.g., $(u(x_{1}),...,u(x_{m}))=\mathcal{E}(u)$. Two separate neural networks (usually multilayer perceptrons), the branch net and the trunk net, serve as the approximator and decoder, respectively. The branch net $b_{\psi}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{r}$ parameterized by $\psi$ processes $\left(u(x_{1}),\ldots,u(x_{m})\right)$ to produce a feature embedding $\left(\beta_{1},\beta_{2},\ldots,\beta_{r}\right)$. The trunk net $t_{\phi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{r}$ with parameters $\phi$, takes a continuous coordinate $y=(y_{1},...,y_{d})$ as input and outputs a feature embedding $\left(\tau_{1},\tau_{2},\ldots,\tau_{r}\right)$. The final DeepONet prediction of a function $u$ for a query $y$ is:
(3) $G_{\boldsymbol{\theta}}(u)(y)=\sum_{k=1}^{r}\beta_{k}\tau_{k}=b_{\psi}(\mathcal{E}(u))\cdot t_{\phi}(y),$
where $\cdot$ is the vector dot product and $\boldsymbol{\theta}=(\psi,\phi)$ represents all the trainable parameters in the branch and trunk nets.
Despite DeepONet’s remarkable success across a range of applications in multiphysics simulation [4, 37, 30], inverse design [36], and carbon storage [17], its supervised training process is highly dependent on the availability of training data, which can be costly. Indeed, the generalization error of DeepONets scales with the number of training input-output function pairs [34, 24, 31]. Generating a large number of high-quality training data is expensive or even impractical in some applications. For example, in simulating high Reynolds number ($Re$) turbulent flow [43], accurate numerical simulations require a fine mesh, leading to a computational cost scaling with $Re^{3}$ [22], making the generation of sufficiently large and diverse training datasets prohibitively expensive.
To address the need for costly data acquisition, physics-informed deep operator networks (PI-DeepONet) [53], inspired by physics-informed neural networks (PINNs) [44], have been proposed to learn operators without relying on observed input-output function pairs. Given a dataset of $N_{f}$ input training functions, $N_{r}$ residual points, $N_{I}$ initial points, and $N_{b}$ boundary points: $\mathcal{D}=\left\{\left\{u^{(i)}\right\}_{i=1}^{N_{f}},\bigl{\{}y_{r}^{(j)}\bigl{\}}_{j=1}^{N_{r}},\bigl{\{}y_{I}^{(j)}\bigl{\}}_{j=1}^{N_{I}},\bigl{\{}y_{b}^{(j)}\bigl{\}}_{j=1}^{N_{b}}\right\}$, PI-DeepONets are trained by minimizing an unsupervised physics loss:
(4) $L_{physics}(\boldsymbol{\theta}|\mathcal{D})=L_{residual}(\boldsymbol{\theta}|\mathcal{D})+\lambda_{I}L_{initial}(\boldsymbol{\theta}|\mathcal{D})+\lambda_{b}L_{boundary}(\boldsymbol{\theta}|\mathcal{D}),$
where
(5) $L_{residual}(\boldsymbol{\theta}|\mathcal{D})$ $=\frac{1}{N_{f}N_{r}}\sum_{i=1}^{N_{f}}\sum_{j=1}^{N_{r}}\Big{|}\mathcal{N}(u^{(i)},G_{\boldsymbol{\theta}}(u^{(i)})(y_{r}^{(j)}))\Big{|}^{2},$
$L_{initial}(\boldsymbol{\theta}|\mathcal{D})$ $=\frac{1}{N_{f}N_{I}}\sum_{i=1}^{N_{f}}\sum_{j=1}^{N_{I}}\Big{|}\mathcal{I}(u^{(i)},G_{\boldsymbol{\theta}}(u^{(i)})(y_{I}^{(j)}))\Big{|}^{2},$
$L_{boundary}(\boldsymbol{\theta}|\mathcal{D})$ $=\frac{1}{N_{f}N_{b}}\sum_{i=1}^{N_{f}}\sum_{j=1}^{N_{b}}\Big{|}\mathcal{B}(u^{(i)},G_{\boldsymbol{\theta}}(u^{(i)})(y_{b}^{(j)}))\Big{|}^{2}.$
Here, $\lambda_{I}$ and $\lambda_{b}$ denote the weight coefficients for different loss terms. However, as noted in the original PI-DeepONet paper [53], the training process can be both memory-intensive and time-consuming. Similar to PINNs [44], this inefficiency arises because optimizing the physics loss requires calculating high-order derivatives of the PDE solution with respect to numerous collocation points, typically achieved via reverse-mode automatic differentiation [2]. This process involves backpropagating the physics loss through the unrolled computational graph to update the model parameters. For PI-DeepONet, the inefficiency is even more pronounced, as the physics loss terms (Equation 5) must be evaluated across multiple PDE configurations. Although various works [6, 13, 7] have proposed different methods to improve the training efficiency of PINNs, little research has focused on enhancing the training efficiency of PI-DeepONet. We propose to address this inefficiency through a separation of input variables.
2.3 Separation of Variables
The method of separation of variables seeks solutions to PDEs of the form $s(y)=T(t)Y_{1}(y_{1})\cdots Y_{d}(y_{d})$ for an input point $y=(t,y_{1},\ldots,y_{d})$ and univariate functions $T,Y_{1},\ldots,Y_{d}$. Suppose we have a linear PDE system
(6) $\mathcal{M}[t]s(y)=\mathcal{L}_{1}[y_{1}]s(y)+\cdots+\mathcal{L}_{d}[y_{d}]s(y),$
where $\mathcal{M}[t]=\frac{d}{dt}+h(t)$ is a first order differential operator of $t$, and $\mathcal{L}_{1}[y_{1}],...,\mathcal{L}_{d}[y_{d}]$ are linear second order ordinary differential operators of their respective variables $y_{1},...,y_{d}$ only. Furthermore, assume we are provided Robin boundary conditions in each variable and separable initial condition $s(t=0,y_{1},\ldots,y_{d})=\prod_{n=1}^{d}\phi_{n}(y_{n})$ for functions $\phi_{n}(y_{n})$ that satisfy the boundary conditions. Then, leveraging Sturm-Liouville theory and some massaging, the solution to this problem can be written
(7) $s(y)=s(t,y_{1},\ldots,y_{d})=\sum_{k}A_{k}T^{k}(t)\prod_{n=1}^{d}Y_{n}^{k}(y_{n}),$
where $k$ is a lumped index that counts over infinite eigenfunctions of each $\mathcal{L}_{i}$ operator (potentially with repeats). For example, given $n\in\{1,...,d\}$, $\mathcal{L}_{n}Y_{n}^{k}(y_{n})=\lambda_{n}^{k}Y_{n}^{k}$ for eigenvalue $\lambda_{n}^{k}\in\mathbb{R}$. $T^{k}(t)$ depends on all the eigenvalues $\lambda_{n}^{k}$ corresponding to index $k$. $A_{k}\in\mathbb{R}$ is a coefficient determined by the initial condition. More details can be found in Appendix C. The method of separation of variables applied to a heat equation example can be found in supplemental materials.
One may notice the resemblance between the form of the DeepONet prediction Eq. 3 with Eq. 7, provided $\beta_{k}=A_{k}$ and $\tau_{k}=T^{k}(t)\prod_{i=1}^{d}Y_{i}^{k}(y_{i})$, with appropriately ordered $k$. We leverage this similarity explicitly in the construction of SepONet below.
3 Separable Operator Networks (SepONet)
Inspired by the method of separation of variables Eq. 7 and recent work on separable PINN [7], we propose using separable operator networks (SepONet) to learn basis functions separately for different coordinate axes. SepONet approximates the solution operator of a PDE system by, for given point $y=(y_{1},\ldots,y_{d})$,
(8) $\begin{split}G_{\boldsymbol{\theta}}(u)(y_{1},\ldots,y_{d})&=\sum_{k=1}^{r}\beta_{k}\prod_{n=1}^{d}\tau_{n,k}\\ &=b_{\psi}(\mathcal{E}(u))\cdot\left(t^{1}_{\phi_{1}}(y_{1})\odot t^{2}_{\phi_{2}}(y_{2})\odot\cdots\odot t_{\phi_{d}}^{d}(y_{d})\right),\end{split}$
where $\odot$ is the Hadamard (element-wise) vector product and $\cdot$ is the vector dot product. Here, $\beta_{k}=b_{\psi}(\mathcal{E}(u))_{k}$ is the $k$-th output of the branch net, as in DeepONet. However, unlike DeepONet, which employs a single trunk net that processes each collocation point $y$ individually, SepONet uses $d$ independent trunk nets, $t^{n}_{\phi_{n}}:\mathbb{R}\rightarrow\mathbb{R}^{r}$ for $n=1,\ldots,d$. In particular, $\tau_{n,k}=t_{\phi_{n}}^{n}(y_{n})_{k}$ denotes the $k$-th output of the $n$-th trunk net. Importantly, the parameters of the $n$-th trunk net $\phi_{n}$ are independent of all other trunk net parameter sets. Viewed through the 3-step approximation framework Eq. 2, SepONet and DeepONet have identical encoder and approximator but different decoders.
Equation 8 can be understood as a low-rank approximation of the solution operator by truncating the basis function series (represented by the output shape of the trunk nets) at a maximal number of ranks $r$. SepONet not only enjoys the advantage that basis functions for different variables with potentially different scales can be learned more easily and efficiently, but also allows for fast and efficient training by leveraging forward-mode automatic differentiation, which we will discuss in Section 3.1. Moreover, despite the resemblance between Eq. 8 and separation of variables method for linear PDEs Eq. 7 (discussed below in Section 3.1.3), we find that SepONet can effectively approximate the solutions to nonlinear parametric PDEs. Indeed, we provide a universal approximation theorem for separable operator networks in Theorem 3.1 and provide extensive accuracy scaling numerical experiments for nonlinear and inseparable PDEs in Section 4.
Finally, it is worth noting that if one is only interested in solving deterministic PDEs under a certain configuration (i.e., $u$ is fixed), then the coefficients $\beta_{k}$ are also fixed and can be absorbed by the basis functions. In this case, SepONet will reduce to separable PINN [7], which has been proven to be efficient and accurate in solving single-scenario PDE systems [9, 39].
3.1 SepONet Architecture and Implementation Details
Suppose we are provided a computation domain $K_{1}=[0,1]^{d}$ of dimension $d$, and an input function $u$. Let $N_{r}=N^{d}$ for given integer $N$. To make predictions along $N^{d}$ collocation points, PI-DeepONet must directly sample all $N^{d}$ points: $\{y^{(i)}\}_{i=1}^{N^{d}}=\{y^{(i)}_{1},...,y^{(i)}_{d}\}_{i=1}^{N^{d}}$ where $y^{(i)}\in K_{1}$ for any $i$. By contrast, to compute predictions at $N_{r}=N^{d}$ collocation points, SepONet only needs to sample $N$ points along each coordinate axis, e.g. $\{\{y_{1}^{(i)}\}_{i=1}^{N},...,\{y_{d}^{(i)}\}_{i=1}^{N}\}$, for a total of $dN$ samples.
For shorthand and generality, we will denote the dataset of input points for SepONet as $\mathcal{D}=\{y_{1}^{(:)},\ldots,y_{d}^{(:)}\}$. Each $y_{n}^{(:)}=\{y_{n}^{(i)}\}_{i=1}^{N_{n}}$ represents an array of $N_{n}$ samples along the $n$-th coordinate axis for a total of $\sum_{n=1}^{d}N_{n}$ samples. The initial and boundary points may be separately sampled from $\partial K_{1}$; the number of samples ($N_{I}$ and $N_{b}$) and sampling strategy are equivalent for SepONet and PI-DeepONet.
3.1.1 Forward Pass
The forward pass of SepONet, illustrated for $d=2$ in Fig. 1, follows the formulation Eq. 8 except generalized to the practical setting where predictions along a grid of collocation points are processed in parallel. The formula can be expressed:
(9) $\begin{split}G_{\boldsymbol{\theta}}(u)(y_{1}^{(:)},\ldots,y_{d}^{(:)})&=\sum_{k=1}^{r}\beta_{k}\bigotimes_{n=1}^{d}\tau_{n,k}^{(:)}\\ &=\sum_{k=1}^{r}b_{\psi}(\mathcal{E}(u))_{k}\left(t^{1}_{\phi_{1}}(y_{1}^{(:)})_{k}\otimes t^{2}_{\phi_{2}}(y_{2}^{(:)})_{k}\otimes\cdots\otimes t_{\phi_{d}}^{d}(y_{d}^{(:)})_{k}\right),\end{split}$
where $\otimes$ is the (outer) tensor product, which produces an output predictive array along a meshgrid of $N_{1}\times N_{2}\times\cdots\times N_{d}$ collocation points. Notably, $\tau_{n,k}^{(:)}=t^{n}_{\phi_{n}}(y_{n}^{(:)})_{k}$ represents a vector of $N_{n}$ values produced by the $n$-th trunk net along the $k$-th output mode after feeding all $y_{n}^{(:)}$ points. After taking the outer product along each of $n=1,\ldots,d$ dimensions for all $r$ modes, the modes are sum-reduced with the predictions of the branch net $\beta_{k}=b_{\psi}(\mathcal{E}(u))_{k}$. While not shown here, our implementation also batches over input functions $\{u^{(i)}\}_{i=1}^{N_{f}}$ for $N_{f}$ functions. Thus, for only $N_{f}+N_{1}+\cdots+N_{d}$ inputs, SepONet produces a predictive array with shape $N_{f}\times N_{1}\times\cdots\times N_{d}$.
3.1.2 Model Update
In evaluation of the physics loss Eq. 4, SepONet enables more efficient computation of high-order derivatives in terms of both time and memory use compared to PI-DeepONet by leveraging forward-mode automatic differentiation (AD) [19]. This is fairly evident by the form of Eq. 9. For example, to compute derivatives along all inputs of the $m$-th axis
(10) $\frac{\partial G_{\theta}(u)(y_{1}^{(:)},...,y_{d}^{(:)})}{\partial y_{m}^{(:)}}$ =$\sum_{k=1}^{r}\beta_{k}\left(\bigotimes_{n\not=m}\tau_{n,k}^{(:)}\right)\otimes\left[\frac{\partial\tau_{m,k}}{\partial y_{m}}\right]^{(:)}$
(11) $:\displaystyle\left[\frac{\partial\tau_{m,k}}{\partial y_{m}}\right]^{(:)}$-$\coloneq\left[\frac{\partial\tau_{m,k}^{(1)}}{\partial y_{m}^{(1)}},\ldots,\frac{\partial\tau_{m,k}^{(N_{m})}}{\partial y_{m}^{(N_{m})}}\right]^{T},:$
where $\left[\frac{\partial\tau_{m,k}}{\partial y_{m}}\right]^{(:)}$ is a vector of derivatives of the $m$-th trunk net’s $k$-th basis function evaluated along all inputs to the $m$-th coordinate axis. One may notice that Eq. 11 can be written as a Jacobian-vector product of the “Jacobian” of the $(m,k)$-th trunk net outputs with respect to all $y_{m}^{(:)}$ inputs with a tangent vector of 1’s: $\left[\frac{\partial\tau_{m,k}}{\partial y_{m}}\right]^{(:)}=\frac{\partial\left(\tau_{m,k}^{(1)},\ldots,\tau_{m,k}^{(N_{m})}\right)}{\partial\left(y_{m}^{(1)},\ldots,y_{m}^{(N_{m})}\right)}\mathbf{1}$. This is equivalent to forward-mode AD. Consequently, the derivatives along the $m$-th coordinate axis across the entire grid of predictions can be obtained by pushing forward derivatives of the $m$-th trunk net, and then reusing the outputs of all other $n\not=m$ trunk nets via outer product. By contrast, PI-DeepOnet must compute derivatives $\frac{\partial t_{\phi}(y_{1},...,y_{d})}{\partial y_{m}}$ for each input-output pair individually, where there is no such advantage and it is more prudent to use reverse-mode AD. Fundamentally, the advantage of SepONet for using forward-mode AD can be attributed to the significantly smaller input-output relationship when evaluating along coordinate grids $\mathbb{R}^{N_{1}+\cdots+N_{d}}\rightarrow\mathbb{R}^{N_{1}\cdots N_{d}}$ compared to PI-DeepONet $\mathbb{R}^{N_{1}\cdots N_{d}}\rightarrow\mathbb{R}^{N_{1}\cdots N_{d}}$. For a more detailed explanation of forward- and reverse-mode AD, we refer readers to [7, 38]. Once the physics loss is computed, reverse-mode AD is employed to update the model parameters $\theta=(\psi,\phi_{1},\ldots,\phi_{d})$.
3.1.3 Inference
Once trained, SepONet can efficiently map input functions $u$ to accurate output function predictions $\hat{s}=G_{\theta}(u)$ along large spatiotemporal grids. This is achieved by combining the learned coefficients from the branch net with the basis functions learned by the trunk nets. Indeed, intriguing comparisons can be made between the results of the method of separation of variables Eq. 7 and trained SepONet predictions Eq. 8. For an initial value problem with linear, separable PDE operator treated in Section 2.3, we might expect SepONet to learn initial value function dependent coefficients $b_{\psi}(\mathcal{E}(u))_{k}=A_{k}$ and spatiotemporal basis functions $t^{n}_{\phi_{n}}(y_{n})_{k}=Y_{n}^{k}(y_{n})$ (provided we supply one additional trunk net $t^{0}_{\phi_{0}}(t)_{k}=T^{k}(t)$ for the temporal dimension) for appropriately ordered modes $k$ and sufficiently large $r$. Examples of the learned basis functions for a separable 1D time-dependent heat equation initial value problem example are provided in the Supplemental Material as a function of the number of the trunk net output shape $r$. For a small number of modes $r=1$ or $r=2$, SepONet learns nearly the exact spatiotemporal basis functions obtained by separation of variables. For larger $r$, the SepONet basis functions do not converge to the analytically expected basis functions. Nevertheless, approximation error is observed to improve with $r$, and near perfect accuracy is obtained at large $r$ by comparison to numerical estimates of the analytic solution.
While the form of SepONet predictions Eq. 8 resemble separation of variables, we note that the method of separation of variables typically only applies to linear PDEs with restricted properties. In spite of this, SepONet is capable of accurately approximating arbitrary operator learning problems (including nonlinear PDEs) as guaranteed by a universal approximation property, provided below.
3.2 Universal Approximation Property of SepONet
The universal approximation property of DeepONet has been discussed in [5, 34]. Here we present the universal approximation theorem to show that proposed separable operator networks can also approximate any nonlinear continuous operators that map infinite-dimensional function spaces to others.
Theorem 3.1 (Universal Approximation Theorem for Separable Operator Networks).
Suppose that $\sigma$ is a Tauber-Wiener function, $g$ is a sinusoidal function, $\mathcal{X}$ is a Banach space, $K\subseteq\mathcal{X}$, $K_{1}\subseteq\mathbb{R}^{d_{1}}$ and $K_{2}\subseteq\mathbb{R}^{d_{2}}$ are three compact sets in $\mathcal{X}$, $\mathbb{R}^{d_{1}}$ and $\mathbb{R}^{d_{2}}$, respectively, $\mathcal{U}$ is a compact set in $C\left(K\right)$, $G$ is a nonlinear continuous operator, which maps $\mathcal{U}$ into a compact set $\mathcal{S}\subseteq C\left(K_{1}\times K_{2}\right)$, then for any $\epsilon>0$, there are positive integers $n$, $r$, $m$, constants $c_{i}^{k}$, $\zeta_{k}^{1}$, $\zeta_{k}^{2}$, $\xi_{ij}^{k}$, $\theta_{i}^{k}\in\mathbb{R}$, points $\omega_{k}^{1}\in\mathbb{R}^{d_{1}}$, $\omega_{k}^{2}\in\mathbb{R}^{d_{2}}$, $x_{j}\in K$, $i=1,\ldots,n$, $k=1,\ldots,r$, $j=1,\ldots,m$, such that
(12) $\left|G(u)(y)-\sum_{k=1}^{r}\underbrace{\sum_{i=1}^{n}c_{i}^{k}\sigma\left(\sum_{j=1}^{m}\xi_{ij}^{k}u\left(x_{j}\right)+\theta_{i}^{k}\right)}_{branch}\underbrace{g\left(w_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)}_{trunk_{1}}\underbrace{g\left(w_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)}_{trunk_{2}}\right|<\epsilon$
holds for all $u\in\mathcal{U}$, $y=(y_{1},y_{2})\in K_{1}\times K_{2}$.
Proof 3.2.
The proof can be found in Section A.2.
Remark 3.3.
The definition of the Tauber-Wiener function is given in Definition A.1. It is worth noting that many common-used activations, such as ReLU, GELU and Tanh, are Tauber-Wiener functions.
Remark 3.4.
Here we show the approximation property of a separable operator network with two trunk nets, by repeatedly applying trigonometric angle addition formula, it is trivial to separate $y$ as $(y_{1},y_{2},\ldots,y_{d})\in K_{1}\times K_{2}\times\ldots\times K_{d}$ and extend Eq. 12 to $d$ trunk nets.
Remark 3.5.
In our assumptions, we restrict the activation function for the trunk nets to be sinusoidal. This choice is motivated by the natural suitability of sinusoidal functions for constructing basis functions [51] and their empirical effectiveness in solving PDEs [50]. However, it would be interesting to explore whether Theorem 3.1 still holds when $g$ is a more general activation function, such as a Tauber-Wiener function. We will leave this investigation for future work.
4 Numerical Results
This section presents comprehensive numerical studies demonstrating the expressive power and effectiveness of SepONet compared to PI-DeepONet on various time-dependent PDEs: diffusion-reaction, advection, Burgers’, and (2+1)-dimensional nonlinear diffusion equations. Both models were trained by optimizing the physics loss (Equation 4) on a dataset $\mathcal{D}$ consisting of input functions, residual points, initial points, and boundary points.
We set the number of residual points $N_{r}=N^{d}=N_{c}$, for problem dimension $d$ and integer $N$. $N_{c}$ will be referred to as the number of training points. The number of initial and boundary points per axis is set to $N_{I}=N_{b}=N=\sqrt[d]{N_{c}}$.
We evaluate both models by varying the number of input functions ($N_{f}$) and training points ($N_{c}$) across four key perspectives: test accuracy, GPU memory usage, training time, and extreme-scale learning capabilities. The main results are illustrated in Fig. 2 and Fig. 3, with complete test results reported in Section B.3. PDE definitions, training details, and problem-specific parameters are provided in Section B.1 and Section B.2.
4.1 Test accuracy
Both PI-DeepONet and SepONet demonstrate improved accuracy when increasing either the number of training points ($N_{c}$) or the number of input functions ($N_{f}$), while fixing the other parameter. This trend is consistent across all four equations tested.
For instance, in the case of the diffusion-reaction equation, when fixing $N_{f}=100$ and varying $N_{c}$ from $8^{2}$ to $128^{2}$, the relative $\ell_{2}$ error of PI-DeepONet decreases from 1.39% to 0.73%, while SepONet’s error reduces from 1.49% to 0.62%. Conversely, when fixing $N_{c}=128^{2}$ and varying $N_{f}$ from 5 to 100, PI-DeepONet’s error drops from 34.54% to 0.73%, and SepONet’s reduces from 22.40% to 0.62%.
4.2 GPU memory usage
While both models show improved accuracy with increasing $N_{c}$ or $N_{f}$, their memory usage patterns differ significantly. This divergence is particularly evident in the case of the advection equation.
When fixing $N_{f}=100$ and varying $N_{c}$, PI-DeepONet exhibits a steep increase in GPU memory consumption, rising from 0.967 GB at $N_{c}=8^{2}$ to 59.806 GB at $N_{c}=128^{2}$. In contrast, SepONet maintains a relatively constant and low memory footprint, ranging between 0.713 GB and 0.719 GB across the same range of $N_{c}$. Similarly, when fixing $N_{c}=128^{2}$ and varying $N_{f}$ from 5 to 100, PI-DeepONet’s memory usage escalates from 3.021 GB to 59.806 GB. SepONet, however, maintains a stable memory usage throughout this range.
4.3 Training time
The training time scaling exhibits a pattern similar to memory usage, as demonstrated by the advection equation example. As $N_{c}$ increases from $8^{2}$ to $128^{2}$ with fixed $N_{f}=100$, PI-DeepONet’s training time increases significantly from 0.0787 to 8.231 hours (2.361 to 246.93 ms per iteration). In contrast, SepONet maintains relatively stable training times, ranging from 0.0730 to 0.0843 hours (2.19 to 2.529 ms per iteration) over the same $N_{c}$ range. Similarly, when varying $N_{f}$ from 5 to 100 with fixed $N_{c}=128^{2}$, PI-DeepONet’s training time increases from 0.3997 to 8.231 hours (11.991 to 246.93 ms per iteration). SepONet, however, keeps training times between 0.0730 and 0.0754 hours. These results demonstrate SepONet’s superior scalability in terms of training time. The ability to maintain near-constant training times across a wide range of problem sizes is a significant advantage, particularly for large-scale applications where computational efficiency is crucial.
4.4 Extreme-scale learning
The Burgers’ and nonlinear diffusion equations highlight SepONet’s capabilities in extreme-scale learning scenarios.
For the Burgers’ equation, PI-DeepONet encounters memory limitations at larger scales. As seen in Fig. 2(c), PI-DeepONet can only compute up to $N_{c}=64^{2}$, achieving a relative $\ell_{2}$ error of 13.72%. In contrast, SepONet continues to improve, reaching a 7.51% error at $N_{c}=128^{2}$. The nonlinear diffusion equation further emphasizes this difference. In Fig. 3(d), PI-DeepONet results are entirely unavailable due to out-of-memory issues. SepONet, however, efficiently handles this complex problem, achieving a relative $\ell_{2}$ error of 6.44% with $N_{c}=128^{3}$ and $N_{f}=100$. Table 1 demonstrates SepONet’s ability to tackle even larger scales for the Burgers’ equation. It achieves a relative $\ell_{2}$ error as low as 4.12% with $N_{c}=512^{2}$ and $N_{f}=800$, while maintaining reasonable memory usage (10.485 GB) and training time (0.478 hours). These results underscore SepONet’s capability to handle extreme-scale learning problems beyond the reach of PI-DeepONet due to computational constraints.
Figure 2: Performance comparison of PI-DeepONet and SepONet with varying number of training points ($N_{c}$) and fixed number of input functions ($N_{f}=100$). Results show test accuracy, GPU memory usage, and training time for four PDEs. As $N_{c}$ increases, both models demonstrate improved accuracy, but PI-DeepONet exhibits significant increases in training time and memory usage, while SepONet maintains better computational efficiency.
Figure 3: Performance comparison of PI-DeepONet and SepONet with increasing number of input functions ($N_{f}$) and fixed number of training points ($N_{c}=128^{d}$, where $d$ is the problem dimension). Both models show improved accuracy with increasing $N_{f}$, but PI-DeepONet’s computational resources scale poorly compared to SepONet’s more efficient scaling. Note: PI-DeepONet results for the (2+1)-dimensional diffusion equation are unavailable due to memory constraints.
Table 1: Additional SepONet results for Burgers’ equation, demonstrating that larger $N_{c}$ and $N_{f}$ can be used to enhance accuracy with minimal cost increase.
Metrics \ $N_{c}$ $N_{f}$ $128^{2}$ 400 $128^{2}$ 800 $256^{2}$ 400 $256^{2}$ 800 $512^{2}$ 400 $512^{2}$ 800
Relative $\ell_{2}$ error (%) 6.60 6.21 5.68 4.46 5.38 4.12
Memory (GB) 0.966 1.466 2.466 4.466 5.593 10.485
Training time (hours) 0.0771 0.0957 0.1238 0.1717 0.2751 0.478
5 Discussion
The field of operator learning faces a critical dilemma. Deep Operator Networks (DeepONets) offer fast training but require extensive data generation, which can be prohibitively expensive or impractical for complex systems. Physics-informed DeepONets (PI-DeepONets) relax the data requirement but at the cost of resource-intensive training processes. This creates a challenging trade-off: balancing resource allocation either before training (data generation) or during training (computational resources).
Our proposed Separable Operator Networks (SepONet) effectively address both of these concerns. Inspired by the separation of variables technique typically used for linear PDEs, SepONet constructs its own basis functions to approximate a nonlinear operator. SepONet’s expressive power is guaranteed by the universal approximation property (Theorem 3.1), ensuring it can approximate any nonlinear continuous operator with arbitrary accuracy. Our numerical results corroborate this theoretical guarantee, demonstrating SepONet’s ability to handle complex, nonlinear systems efficiently.
By leveraging independent trunk networks for different variables, SepONet enables an efficient implementation via forward-mode automatic differentiation (AD). This approach achieves remarkable efficiency in both data utilization and computational resources. SepONet is trained solely by optimizing the physics loss, eliminating the need for expensive simulations to generate ground truth PDE solutions. In terms of computational resources, SepONet maintains stable GPU memory usage and training time, even with increasing training data and network size, in contrast to PI-DeepONet’s dramatic resource consumption increases under similar scaling.
However, SepONet has certain limitations that warrant further research and development. The mesh grid structure of SepONet’s solution, while enabling efficient training through forward-mode AD, may limit its flexibility in handling PDEs with irregular geometries. Addressing this limitation could involve developing adaptations or hybrid approaches that accommodate more complex spatial domains [26, 48, 10], potentially expanding SepONet’s applicability to a broader range of physical problems.
Additionally, while the linear decoder allows for an efficient SepONet implementation, a very large number of basis functions may be needed for accurate linear representation in some problems [47]. Developing a nonlinear decoder version of SepONet will be useful to balance accuracy and efficient training. Moreover, implementing an adaptive weighting strategy [52, 54, 55] for different loss terms in the physics loss function, instead of using predefined fixed weights, could lead to improved accuracy and faster convergence.
Finally, empirical observations suggest that training accuracy and robustness improve with an increase in input training functions. However, the neural scaling laws in physics-informed operator learning remain unexplored, presenting an intriguing theoretical challenge for future investigation.
Acknowledgments
We would like to thank Wolfger Peelaers for the valuable discussions and insightful comments.
Appendix A Universal Approximation Theorem for Separable Operator Networks
Here we present the universal approximation theorem for the proposed separable operator networks, originally written in Theorem 3.1 and repeated below in Theorem A.14. We begin by reviewing established theoretical results on approximating continuous functions and functionals. Following this review, we introduce the preliminary lemmas and proofs necessary for understanding Theorem A.14. We refer our readers to [5, 57] for detailed proofs of Theorem A.3, Theorem A.4, Theorem A.5. Main notations are listed in Table 2.
A.1 Preliminaries and Auxiliary Results
Definition A.1 (Tauber-Wiener (TW)).
If a function $g:\mathbb{R}\rightarrow\mathbb{R}$ (continuous or discontinuous) satisfies that all the linear combinations $\sum_{i=1}^{N}c_{i}g\left(\lambda_{i}x+\theta_{i}\right)$, $\lambda_{i}\in\mathbb{R}$, $\theta_{i}\in\mathbb{R}$, $c_{i}\in\mathbb{R}$, $i=1,2,\ldots,N$, are dense in every $C[a,b]$, then $g$ is called a Tauber-Wiener (TW) function.
{remark_appendix}
[Density in $C[a,b]$] A set of functions is said to be dense in $C[a,b]$ if every function in the space of continuous functions on the interval $[a,b]$ can be approximated arbitrarily closely by functions from the set.
Definition A.2 (Compact Set).
Suppose that $X$ is a Banach space, $V\subseteq X$ is called a compact set in $X$, if for every sequence $\left\{x_{n}\right\}_{n=1}^{\infty}$ with all $x_{n}\in V$, there is a subsequence $\left\{x_{n_{k}}\right\}$, which converges to some element $x\in V$.
Theorem A.3 ([5]).
Suppose that $K$ is a compact set in $\mathbb{R}^{n}$, $\mathcal{S}$ is a compact set in $C(K)$, $g\in(\mathrm{TW})$, then for any $\epsilon>0$, there exist a positive integer $N$, real numbers $\theta_{i}$, vectors $\omega_{i}\in\mathbb{R}^{n}$, $i=1,\ldots,N$, which are independent of $f\in C(K)$ and constants $c_{i}(f)$, $i=1,\ldots,N$ depending on $f$, such that
(13) $\left|f(x)-\sum_{i=1}^{N}c_{i}(f)g\left(\omega_{i}\cdot x+\theta_{i}\right)\right|<\epsilon$
holds for all $x\in K$ and $f\in\mathcal{S}$. Moreover, each $c_{i}(f)$ is a linear continuous functional defined on $\mathcal{S}$.
Theorem A.4 ([5]).
Suppose that $\sigma\in(\mathrm{TW}),X$ is a Banach Space, $K\subseteq X$ is a compact set, $\mathcal{U}$ is a compact set in $C(K)$, $f$ is a continuous functional defined on $\mathcal{U}$, then for any $\epsilon>0$, there are positive integers $N$, $m$ points $x_{1},\ldots,x_{m}\in K$, and real constants $c_{i}$, $\theta_{i}$, $\xi_{ij}$, $i=1,\ldots,N$, $j=1,\ldots,m$, such that
(14) $\left|f(u)-\sum_{i=1}^{N}c_{i}\sigma\left(\sum_{j=1}^{m}\xi_{ij}u\left(x_{j}\right)+\theta_{i}\right)\right|<\epsilon$
holds for all $u\in\mathcal{U}$.
Theorem A.5 (Weierstrass Approximation Theorem [57]).
Suppose $f\in C[a,b],$ then for every $\epsilon>0,$ there exists a polynomial p such that for all $x$ in $[a,b]$, we have $|f(x)-p(x)|<\epsilon.$
Table 2: Notations and Symbols
$X$ some Banach space with norm $\|\cdot\|_{X}$
$\mathbb{R}^{d}$ Euclidean space of dimension $d$
$K$ some compact set in a Banach space
$C(K)$ Banach space of all continuous functions defined on $K$, with norm $\|f\|_{C(K)}=\max_{x\in K}|f(x)|$
$\tilde{C}[a,b]$ the space of functions in $C[a,b]$ satisfying $f(a)=f(b)$
$V$ some compact set in $C(K)$
$u(x)$ some input function
$\mathcal{U}$ the space of input functions
$G$ some continuous operator
$G(u)(y)$ or $s(y)$ some output function that is mapped from the corresponding input function $u$ by the operator $G$
$\mathcal{S}$ the space of output functions
(TW) all the Tauber-Wiener functions
$\sigma$ and $g$ activation function for branch net and trunk nets in Theorem A.14
$\left\{x_{1},x_{2},\ldots,x_{m}\right\}$ $m$ sensor points for identifying input function $u$
$r$ rank of some deep operator network or separable operator network
$n,m$ operator network size hyperparameters in Theorem A.14
Corollary A.6.
Trigonometric polynomials are dense in the space of continuous and periodic functions $\tilde{C}[0,2\pi]:=\{f\in C[0,2\pi]~{}|~{}f(0)=f(2\pi)\}$.
Proof A.7.
For any $\tilde{f}\in\tilde{C}[0,2\pi]$, extend it to a $2\pi$-periodic and continuous function $f$ defined on $\mathbb{R}$. It suffices to show that there exists a trigonometric polynomial that approximates $f$ within any $\epsilon>0$. We construct the continuous even functions of $2\pi$ period $g$ and $h$ as:
(15) $g(\theta)=\frac{f(\theta)+f(-\theta)}{2}\quad\text{and}\quad h(\theta)=\frac{f(\theta)-f(-\theta)}{2}\sin(\theta).$
Let $\phi(x)=g(\arccos x)$ and $\psi(x)=h(\arccos x)$. Since $\phi,\psi$ are continuous functions on $[-1,1]$, by the Weierstrass Approximation Theorem Theorem A.5, for any $\epsilon>0$, there exist polynomials $p$ and $q$ such that
(16) $|\phi(x)-p(x)|<\frac{\epsilon}{4}\quad\text{and}\quad|\psi(x)-q(x)|<\frac{\epsilon}{4}$
holds for all $x\in[-1,1]$. Let $x=\cos\theta$, it follows that
(17) $|g(\theta)-p(\cos\theta)|<\frac{\epsilon}{4}\quad\text{and}\quad|h(\theta)-q(\cos\theta)|<\frac{\epsilon}{4}$
for $\theta\in[0,\pi]$. Because $g$, $h$ and $cosine$ are even and $2\pi$-periodic, Eq. 17 holds for all $\theta\in\mathbb{R}$. From the definitions of $g$ and $h$, and the fact $\left|\sin\theta\right|\leq 1$, $\left|\sin^{2}\theta\right|\leq 1$, we have
(18) $\left|\frac{f(\theta)+f(-\theta)}{2}\sin^{2}\theta-p(\cos\theta)\sin^{2}\theta\right|<\frac{\epsilon}{4}\quad\text{and}\quad\left|\frac{f(\theta)-f(-\theta)}{2}\sin^{2}\theta-q(\cos\theta)\sin\theta\right|<\frac{\epsilon}{4}.$
Using the triangle inequality, we obtain
(19) $\left|f(\theta)\sin^{2}\theta-\left[p(\cos\theta)\sin^{2}\theta+q(\cos\theta)\sin\theta\right]\right|<\frac{\epsilon}{2}.$
Applying the same analysis to
(20) $\tilde{g}(\theta)=\frac{f(\theta+\frac{\pi}{2})+f(-\theta+\frac{\pi}{2})}{2}\quad\text{and}\quad\tilde{h}(\theta)=\frac{f(\theta+\frac{\pi}{2})-f(-\theta+\frac{\pi}{2})}{2}\sin(\theta),$
we can find polynomials $r$ and $s$ such that
(21) $\left|f\left(\theta+\frac{\pi}{2}\right)\sin^{2}\theta-\left[r(\cos\theta)\sin^{2}\theta+s(\cos\theta)\sin\theta\right]\right|<\frac{\epsilon}{2}$
holds for all $\theta$. Substituting $\theta$ with $\theta-\frac{\pi}{2}$ gives
(22) $\left|f(\theta)\cos^{2}\theta-\left[r(\sin\theta)\cos^{2}\theta-s(\sin\theta)\cos\theta\right]\right|<\frac{\epsilon}{2}.$
By the triangle inequality, combining Eq. 22 and Eq. 19 gives
(23) $\left|f(\theta)-\left[r(\sin\theta)\cos^{2}\theta-s(\sin\theta)\cos\theta+p(\cos\theta)\sin^{2}\theta+q(\cos\theta)\sin\theta\right]\right|<\epsilon$
holds for all $\theta$. Thus, the trigonometric polynomial
(24) $r(\sin\theta)\cos^{2}\theta-s(\sin\theta)\cos\theta+p(\cos\theta)\sin^{2}\theta+q(\cos\theta)\sin\theta$
is an $\epsilon$-approximation to $f$.
{remark_appendix}
If $p(x)$ is a polynomial, it is easy to verify that $p(\cos\theta)$ is a trigonometric polynomial due to the fact $\cos^{n}\theta=\sum_{k=0}^{n}\frac{\binom{n}{k}}{2^{n}}\cos\left((n-2k)\theta\right)$. Prior to proving Theorem 3.1, we need to establish the following lemmas.
Lemma A.8.
Sine is a Tauber-Wiener function.
Proof A.9.
Assuming the interval to be $[0,\pi]$ first. For every continuous function $f$ on $[0,\pi]$ and any $\epsilon>0$, we can extend $f$ to a continuous function $F$ on $[0,2\pi]$ so that $F(x)=f(x)$ on $[0,\pi]$ and $F(2\pi)=F(0)$. By Corollary A.6, there exists a trigonometric polynomial
(25) $p(x)=a_{0}+\sum_{n=1}^{N}a_{n}\cos(nx)+b_{n}\sin(nx)$
such that
(26) $\sup_{x\in[0,\pi]}|f(x)-p(x)|\leq\sup_{x\in[0,2\pi]}|F(x)-p(x)|<\epsilon.$
Let $c_{0}=a_{0}$, $\lambda_{0}=0$, $\theta_{0}=\frac{\pi}{2}$, $c_{2n-1}=b_{n}$, $\lambda_{2n-1}=n$, $\theta_{2n-1}=0$, $c_{2n}=a_{n}$, $\lambda_{2n}=n$, $\theta_{2n}=\frac{\pi}{2}$, for $n=1,2,\ldots,N$, $p(x)$ is redefined as
(27) $p(x)=\sum_{i=0}^{2N}c_{i}\sin\left(\lambda_{i}x+\theta_{i}\right).$
Thus we have
(28) $\left|f(x)-\sum_{i=0}^{2N}c_{i}\sin\left(\lambda_{i}x+\theta_{i}\right)\right|<\epsilon$
for $x\in[0,\pi]$. Now consider a continuous function $g$ on $[a,b]$, define $f(x)\in C[0,\pi]:=g\left(\frac{b-a}{\pi}x+a\right)$, then by Eq. 28, we have
(29) $\left|g(x)-\sum_{i=0}^{2N}c_{i}\sin\left(\frac{\pi\lambda_{i}}{b-a}x-\frac{\pi\lambda_{i}a}{b-a}+\theta_{i}\right)\right|<\epsilon$
holds for all $x\in[a,b]$. Therefore, it follows that for any continuous function $g$ on $[a,b]$ and any $\epsilon>0$, we can approximate $g$ within $\epsilon$ by choosing $N$ sufficiently large and adjusting $c_{i}$, $\lambda_{i}$, $\theta_{i}$ accordingly. Hence, the set of all such linear combinations of $\sin(x)$ is dense in $C[a,b]$, confirming that $\sin(x)$ is a Tauber-Wiener function.
{remark_appendix}
It is straightforward to conclude that all sinusoidal functions are Tauber-Wiener functions.
Lemma A.10.
Suppose that $V_{1}\subseteq X_{1}$, $V_{2}\subseteq X_{2}$ are two compact sets in Banach spaces $X_{1}$ and $X_{2}$, respectively, then their Cartesian product $V_{1}\times V_{2}$ is also compact.
Proof A.11.
For every sequence $\left\{x_{n}^{1},x_{n}^{2}\right\}$ in $V_{1}\times V_{2}$, since $V_{1}$ is compact, $\left\{x_{n}^{1}\right\}$ has a subsequence $\left\{x_{n_{k}}^{1}\right\}$ that converges to some element $x^{1}\in V_{1}$. As well, since $V_{2}$ is compact, there exists a subsequence $\left\{x_{n_{k}}^{2}\right\}$ that converges to $x^{2}\in V_{2}$. It follows that $\left\{x_{n}^{1},x_{n}^{2}\right\}$ converges to $\left(x^{1},x^{2}\right)\in V_{1}\times V_{2}$, thus $V_{1}\times V_{2}$ is compact.
Lemma A.12.
Suppose that $X$ is a Banach space, $K_{1}\subseteq X_{1}$, $K_{2}\subseteq X_{2}$ are two compact sets in $X_{1}$ and $X_{2}$, respectively. $\mathcal{U}$ is a compact set in $C(K_{1})$, then the range $G(\mathcal{U})$ of the continuous operator $G$ from $\mathcal{U}$ to $C(K_{2})$ is compact in $C(K_{2})$.
Proof A.13.
For every sequence $\{f_{n}\}$ in $\mathcal{U}$, since $\mathcal{U}$ is compact, there exists a subsequence $\{f_{n_{k}}\}$ that converges to some function $f\in\mathcal{U}$. Since $G$ is continuous, the convergence $f_{n_{k}}\to f$ in $C(K_{1})$ implies
(30) $G(f_{n_{k}})\to G(f)\quad\text{in}\quad C(K_{2}).$
Thus, for every sequence $\{G(f_{n})\}$ in $G(\mathcal{U})$, there exists a subsequence $\{G(f_{n_{k}})\}$ that converges to $G(f)\in G(\mathcal{U})$. Therefore, the range $G(\mathcal{U})$ of the continuous operator $G$ is compact in $C(K_{2})$.
A.2 Universal Approximation Theorem for SepONet
Theorem A.14 (Universal Approximation Theorem for Separable Operator Networks).
Suppose that $\sigma\in(\mathrm{TW})$, $g$ is a sinusoidal function, $X$ is a Banach Space, $K\subseteq X$, $K_{1}\subseteq\mathbb{R}^{d_{1}}$ and $K_{2}\subseteq\mathbb{R}^{d_{2}}$ are three compact sets in $X$, $\mathbb{R}^{d_{1}}$ and $\mathbb{R}^{d_{2}}$, respectively, $\mathcal{U}$ is a compact set in $C\left(K\right)$, $G$ is a nonlinear continuous operator, which maps $\mathcal{U}$ into a compact set $\mathcal{S}\subseteq C\left(K_{1}\times K_{2}\right)$, then for any $\epsilon>0$, there are positive integers $n$, $r$, $m$, constants $c_{i}^{k}$, $\zeta_{k}^{1}$, $\zeta_{k}^{2}$, $\xi_{ij}^{k}$, $\theta_{i}^{k}\in\mathbb{R}$, points $\omega_{k}^{1}\in\mathbb{R}^{d_{1}}$, $\omega_{k}^{2}\in\mathbb{R}^{d_{2}}$, $x_{j}\in K_{1}$, $i=1,\ldots,n$, $k=1,\ldots,r$, $j=1,\ldots,m$, such that
(31) $\left|G(u)(y)-\sum_{k=1}^{r}\sum_{i=1}^{n}c_{i}^{k}\sigma\left(\sum_{j=1}^{m}\xi_{ij}^{k}u\left(x_{j}\right)+\theta_{i}^{k}\right)g\left(w_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)g\left(w_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)\right|<\epsilon$
holds for all $u\in\mathcal{U}$, $y=(y_{1},y_{2})\in K_{1}\times K_{2}$.
Proof A.15.
Without loss of generality, we can assume that $g$ is sine function, by Lemma A.8, we have $g\in(TW)$; From the assumption that $K_{1}$ and $K_{2}$ are compact, by Lemma A.10, $K_{1}\times K_{2}$ is compact; Since $G$ is a continuous operator that maps $\mathcal{U}$ into $C(K_{1}\times K_{2})$, it follows that the range $G(\mathcal{U})=\left\{G(u):u\in\mathcal{U}\right\}$ is compact in $C(K_{1}\times K_{2})$ due to Lemma A.12; Thus by Theorem A.3, for any $\epsilon>0$, there exists a positive integer $N$, real numbers $c_{k}(G(u))$ and $\zeta_{k}$, vectors $\omega_{k}\in R^{d_{1}+d_{2}}$, $k=1,\ldots,N$, such that
(32) $\left|G(u)(y)-\sum_{k=1}^{N}c_{k}(G(u))g\left(\omega_{k}\cdot y+\zeta_{k}\right)\right|<\frac{\epsilon}{2}$
holds for all $y\in K_{1}\times K_{2}$ and $u\in C(K)$. Let $(\omega_{k}^{1},\omega_{k}^{2})=\omega_{k}$, where $\omega_{k}^{1}\in\mathbb{R}^{d_{1}}$ and $\omega_{k}^{2}\in\mathbb{R}^{d_{2}}$. Utilizing the trigonometric angle addition formula, we have
(33) $g\left(\omega_{k}\cdot y+\zeta_{k}\right)=g\left(\omega_{k}^{1}\cdot y_{1}+\zeta_{k}\right)g\left(\omega_{k}^{2}\cdot y_{2}+\frac{\pi}{2}\right)+g\left(\omega_{k}^{1}\cdot y_{1}+\zeta_{k}+\frac{\pi}{2}\right)g\left(\omega_{k}^{2}\cdot y_{2}\right).$
Let $r=2N$, $c_{N+k}(G(u))=c_{k}(G(u))$, $\omega_{N+k}^{1}=\omega_{k}^{1}$, $\omega_{N+k}^{2}=\omega_{k}^{2}$, $\zeta_{k}^{1}=\zeta_{k}$, $\zeta_{k}^{2}=\frac{\pi}{2}$ for $k=1,\ldots,N$, and let $\zeta_{k}^{1}=\zeta_{k}+\frac{\pi}{2}$, $\zeta_{k}^{2}=0$ for $k=N+1,\ldots,r$, Eq. 32 can be expressed as:
(34) $\left|G(u)(y)-\sum_{k=1}^{r}c_{k}(G(u))g\left(\omega_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)g\left(\omega_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)\right|<\frac{\epsilon}{2}.$
Since $G$ is a continuous operator, according to the last proposition of Theorem A.3, we conclude that for each $k=1,\ldots,2N$, $c_{k}(G(u))$ is a continuous functional defined on $\mathcal{U}$. Repeatedly applying Theorem A.4, for each $k=1,\ldots,2N$, $c_{k}(G(u))$, we can find positive integers $n_{k}$,$m_{k}$, constants $c_{i}^{k}$, $\xi_{ij}^{k}$, $\theta_{i}^{k}\in R$ and $x_{j}\in K_{1}$, $i=1,\ldots,n_{k}$, $j=1,\ldots,m_{k}$, such that
(35) $\left|c_{k}(G(u))-\sum_{i=1}^{n_{k}}c_{i}^{k}\sigma\left(\sum_{j=1}^{m_{k}}\xi_{ij}^{k}u(x_{j})+\theta_{i}^{k}\right)\right|<\frac{\epsilon}{2L}$
holds for all $k=1,\ldots,r$ and $u\in\mathcal{U}$, where
(36) $L=\sum_{k=1}^{r}\sup_{y_{1}\in K_{2},y_{2}\in K_{3}}\left|g\left(\omega_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)g\left(\omega_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)\right|.$
Substituting Eq. 35 into Eq. 34, we obtain that
(37) $\left|G(u)(y)-\sum_{k=1}^{r}\sum_{i=1}^{n_{k}}c_{i}^{k}\sigma\left(\sum_{j=1}^{m_{k}}\xi_{ij}^{k}u\left(x_{j}\right)+\theta_{i}^{k}\right)g\left(w_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)g\left(w_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)\right|<\epsilon$
holds for all $u\in\mathcal{U}$, $y_{1}\in K_{1}$ and $y_{2}\in K_{2}$. Let $n=\max_{k}{n_{k}}$, $m=\max_{k}{m_{k}}$. For all $n_{k}<i\leq n$, let $c_{i}^{k}=0$. For all $m_{k}<j\leq m$, let $\xi_{ij}^{k}=0$. Then Eq. 37 can be rewritten as:
(38) $\left|G(u)(y)-\sum_{k=1}^{r}\sum_{i=1}^{n}c_{i}^{k}\sigma\left(\sum_{j=1}^{m}\xi_{ij}^{k}u\left(x_{j}\right)+\theta_{i}^{k}\right)g\left(w_{k}^{1}\cdot y_{1}+\zeta_{k}^{1}\right)g\left(w_{k}^{2}\cdot y_{2}+\zeta_{k}^{2}\right)\right|<\epsilon,$
which holds for all $u\in\mathcal{U}$, $y_{1}\in K_{1}$ and $y_{2}\in K_{2}$. This completes the proof of Theorem 3.1.
Appendix B PDE Problem Definitions, Training details, and Complete Test Results
B.1 PDE Problem Definitions
All PDE test problems exhibited in Section 4 are described in the subsections below.
B.1.1 Diffusion-Reaction Systems
Consider the following nonlinear diffusion-reaction system with a source term $u(x)$:
(39) $\frac{\partial s}{\partial t}=D\frac{\partial^{2}s}{\partial x^{2}}+ks^{2}+u(x),\quad(x,t)\in(0,1)\times(0,1]$
with zero initial and boundary conditions, where $D=0.01$ is the diffusion coefficient and $k=0.01$ is the reaction rate. The input training source terms are sampled from a mean-zero Gaussian random field (GRF) [46] with a length scale 0.2. To generate the test dataset, we sample 100 different source terms from the same GRF and apply a second-order implicit finite difference method [16] to obtain the reference solutions on a uniform $128\times 128$ grid.
B.1.2 Advection equation
Consider the following linear advection equation parameterized by the variable coefficients $u(x)$:
(40) $\frac{\partial s}{\partial t}+u(x)\frac{\partial s}{\partial x}=0,\quad(x,t)\in(0,1)\times(0,1)$
with the initial and boundary condition
(41) $s(x,0)=\sin(\pi x),\quad x\in(0,1),$
$s(0,t)=\sin\left(\frac{\pi}{2}t\right),\quad t\in(0,1).$
The input training variable coefficients are strictly positive by defining $u(x)=v(x)-\min_{x}v(x)+1$, where $v$ is sampled from a GRF with length scale 0.2. To create the test dataset, we generate 100 new coefficients in the same manner that are not used in training and apply the Lax–Wendroff scheme [16] to solve the advection equation on a uniform $128\times 128$ grid.
B.1.3 Burgers’ Equation
Consider the nonlinear Burgers’ equation:
(42) $\frac{\partial s}{\partial t}+s\frac{\partial s}{\partial x}-\nu\frac{\partial^{2}s}{\partial x^{2}}=0,\quad(x,t)\in(0,1)\times(0,1],$
$s(x,0)=u(x),\quad x\in(0,1)$
with periodic boundary conditions:
(43) $s(0,t)=s(1,t),$
$\frac{\partial s}{\partial x}(0,t)=\frac{\partial s}{\partial x}(1,t),$
where $t\in(0,1)$ and $\nu=0.01$ is the viscosity. The input training initial conditions are sampled from a GRF $\sim\mathcal{N}\left(0,25^{2}\left(-\Delta+5^{2}I\right)^{-4}\right)$ using the Chebfun package [8], satisfying the periodic boundary conditions. Synthetic test dataset consists of 100 unseen initial functions and their corresponding solutions, which are generated from the same GRF and are solved by spectral method on a $101\times 101$ uniform grid using the spinOp library [41], respectively.
B.1.4 2D Nonlinear diffusion equation
Consider the 2D nonlinear diffusion equation which was used in [7]:
(44) $\frac{\partial s}{\partial t}=\alpha\nabla\cdot\left(s\nabla s\right),\quad(\boldsymbol{x},t)\in\Omega\times[0,1],$
$s(\boldsymbol{x},0)=u(\boldsymbol{x}),\quad\boldsymbol{x}\in\Omega,$
$s(\boldsymbol{x},t)=0,\quad(\boldsymbol{x},t)\in\partial\Omega\times[0,1],$
where $\boldsymbol{x}=(x,y)$ denotes 2D spatial variables, $\Omega=[0,1]^{2}$ is the spatial domain and $\alpha=0.05$ is the diffusivity. The input training initial conditions are generated as a sum of Gaussian functions, parameterized as:
(45) $u(x,y)=\sum_{i=1}^{3}A_{i}\exp[-w_{i}\{(x-x_{i})^{2}+(y-y_{i})^{2}\}],$
where $A_{i}\sim\mathcal{U}(0.2,0.5)$ are amplitudes, $w_{i}\sim\mathcal{U}(10,20)$ are width parameters, and $(x_{i},y_{i})\sim\mathcal{U}(-0.5,0.5)^{2}$ are center coordinates. We also generate 100 unseen test initial conditions using this method. The nonlinear diffusion equation is then solved using explicit Adams method to obtain reference solutions on a uniform $101\times 101$ spatial grid with 101 time points.
B.2 Training details and hyperparameters
Both PI-DeepONet and SepONet were trained by minimizing the physics loss (Equation 4) using gradient descent with the Adam optimizer [21]. The initial learning rate is $1\times 10^{-3}$ and decays by a factor of 0.9 every 1,000 iterations. Additionally, we resample input training functions and training points (including residual, initial, and boundary points) every 100 iterations.
Across all benchmarks and on both models (SepONet and PI-DeepONet), we apply Tanh activation for the branch net and Sine activation for the trunk net. We note that no extensive hyperparameter tuning was performed for either PI-DeepONet or SepONet. The code in this study is implemented using JAX and Equinox libraries [3, 20], and all training was performed on a single NVIDIA A100 GPU with 80 GB of memory. Training hyperparameters are provided in Table 3. All code accompanying these benchmarks is available at https://github.com/HewlettPackard/separable-operator-networks.
Table 3: Training hyperparameters for different PDE benchmarks
Hyperparameters \ PDEs Diffusion-reaction Advection Burgers’ 2D Nonlinear diffusion
# of sensors 128 128 101 10201 (101 × 101)
Network depth 5 6 7 7
Network width 50 100 100 128
# of training iterations 50k 120k 80k 80k
Weight coefficients ($\lambda_{I}$ / $\lambda_{b}$) 1 / 1 100 / 100 20 / 1 20 / 1
B.3 Complete test results
We report the relative $\ell_{2}$ error, root mean squared error (RMSE), GPU memory usage and total training time as metrics to assess the performance of PI-DeepONet and SepONet. Specifically, the mean and standard deviation of the relative $\ell_{2}$ error and RMSE are calculated over all functions in the test dataset. The complete test results are shown in Table 4 and Table 5.
Table 4: Performance comparison of PI-DeepONet and SepONet with varying number of training points ($N_{c}$) and fixed number of input training functions ($N_{f}=100$). The ’-’ symbol indicates that results are not available due to out-of-memory issues.
Equations Metrics Models $8^{d}$ $16^{d}$ $32^{d}$ $64^{d}$ $128^{d}$
Diffusion-Reaction $d=2$ Relative $\ell_{2}$ error (%) PI-DeepONet 1.39 ± 0.71 1.11 ± 0.59 0.87 ± 0.41 0.83 ± 0.35 0.73 ± 0.34
SepONet 1.49 ± 0.82 0.79 ± 0.35 0.70 ± 0.33 0.62 ± 0.28 0.62 ± 0.26
RMSE (×$\times 10^{-2}$) PI-DeepONet 0.58 ± 0.29 0.46 ± 0.22 0.37 ± 0.20 0.36 ± 0.20 0.32 ± 0.18
SepONet 0.62 ± 0.28 0.35 ± 0.22 0.32 ± 0.23 0.28 ± 0.20 0.29 ± 0.21
Memory (GB) PI-DeepONet 0.729 1.227 3.023 9.175 35.371
SepONet 0.715 0.717 0.715 0.717 0.719
Training time (hours) PI-DeepONet 0.0433 0.0641 0.1497 0.7252 2.8025
SepONet 0.0403 0.0418 0.0430 0.0427 0.0326
Advection $d=2$ Relative $\ell_{2}$ error (%) PI-DeepONet 9.27 ± 1.94 7.55 ± 1.86 6.79 ± 1.84 6.69 ± 1.95 5.72 ± 1.57
SepONet 14.29 ± 2.65 11.96 ± 2.17 6.14 ± 1.58 5.80 ± 1.57 4.99 ± 1.40
RMSE (×$\times 10^{-2}$) PI-DeepONet 5.88 ± 1.34 4.79 ± 1.27 4.31 ± 1.23 4.24 ± 1.29 3.63 ± 1.05
SepONet 9.06 ± 1.88 7.58 ± 1.55 3.90 ± 1.09 3.69 ± 1.07 3.17 ± 0.95
Memory (GB) PI-DeepONet 0.967 1.741 5.103 17.995 59.806
SepONet 0.713 0.715 0.715 0.715 0.719
Training time (hours) PI-DeepONet 0.0787 0.1411 0.4836 2.3987 8.231
SepONet 0.0843 0.0815 0.0844 0.0726 0.0730
Burgers’ $d=2$ Relative $\ell_{2}$ error (%) PI-DeepONet 29.33 ± 3.85 20.31 ± 4.31 14.17 ± 5.25 13.72 ± 5.59 -
SepONet 29.42 ± 3.79 31.53 ± 3.44 28.74 ± 4.11 11.85 ± 4.06 7.51 ± 4.04
RMSE (×$\times 10^{-2}$) PI-DeepONet 4.19 ± 2.79 2.82 ± 1.86 2.23 ± 2.10 2.20 ± 2.13 -
SepONet 4.18 ± 2.74 4.44 ± 2.81 4.11 ± 2.76 1.80 ± 1.60 1.23 ± 1.44
Memory (GB) PI-DeepONet 1.253 2.781 5.087 18.001 -
SepONet 0.603 0.605 0.603 0.605 0.716
Training time (hours) PI-DeepONet 0.1497 0.2375 0.6431 3.2162 -
SepONet 0.0706 0.0719 0.0716 0.0718 0.0605
Nonlinear diffusion $d=3$ Relative $\ell_{2}$ error (%) PI-DeepONet 17.38 ± 5.56 9.90 ± 2.91 - - -
SepONet 16.10 ± 4.46 12.11 ± 3.89 6.81 ± 1.98 6.73 ± 1.96 6.44 ± 1.69
RMSE (×$\times 10^{-2}$) PI-DeepONet 1.86 ± 0.62 1.04 ± 0.23 - - -
SepONet 1.72 ± 0.49 1.29 ± 0.37 0.72 ± 0.19 0.71 ± 0.17 0.68 ± 0.15
Memory (GB) PI-DeepONet 6.993 37.715 - - -
SepONet 3.471 2.897 2.899 2.897 13.139
Training time (hours) PI-DeepONet 0.5836 6.6399 - - -
SepONet 0.1044 0.1069 0.1056 0.1456 0.5575
Table 5: Performance comparison of PI-DeepONet and SepONet with varying number of input training functions ($N_{f}$) and fixed number of training points ($N_{c}=128^{d}$, $d$ indicated by problem instance). The ’-’ symbol indicates that results are not available due to out-of-memory issues.
Equations Metrics Models $5$ $10$ $20$ $50$ $100$
Diffusion-Reaction $d=2$ Relative $\ell_{2}$ error (%) PI-DeepONet 34.54 ± 27.83 4.23 ± 2.52 1.72 ± 1.00 0.91 ± 0.46 0.73 ± 0.34
SepONet 22.40 ± 12.30 3.11 ± 1.89 1.19 ± 0.74 0.73 ± 0.32 0.62 ± 0.26
RMSE (×$\times 10^{-2}$) PI-DeepONet 14.50 ± 9.04 1.75 ± 0.90 0.71 ± 0.37 0.40 ± 0.28 0.32 ± 0.18
SepONet 9.34 ± 5.37 1.36 ± 1.07 0.50 ± 0.29 0.34 ± 0.25 0.29 ± 0.21
Memory (GB) PI-DeepONet 2.767 5.105 9.239 17.951 35.371
SepONet 0.719 0.719 0.717 0.717 0.719
Training time (hours) PI-DeepONet 0.1268 0.2218 0.5864 1.4018 2.8025
SepONet 0.0375 0.0390 0.0370 0.0317 0.0326
Advection $d=2$ Relative $\ell_{2}$ error (%) PI-DeepONet 9.64 ± 2.91 8.77 ± 2.23 7.57 ± 1.98 6.69 ± 1.93 5.72 ± 1.57
SepONet 7.62 ± 2.06 6.59 ± 1.71 5.47 ± 1.57 5.18 ± 1.51 4.99 ± 1.40
RMSE (×$\times 10^{-2}$) PI-DeepONet 6.11 ± 1.90 5.55 ± 1.51 4.80 ± 1.33 4.24 ± 1.28 3.63 ± 1.05
SepONet 4.83 ± 1.38 4.18 ± 1.17 3.47 ± 1.06 3.29 ± 1.02 3.17 ± 0.95
Memory (GB) PI-DeepONet 3.021 5.611 9.707 34.511 59.806
SepONet 0.713 0.715 0.719 0.719 0.719
Training time (hours) PI-DeepONet 0.3997 1.0766 1.9765 4.411 8.231
SepONet 0.0754 0.0715 0.0736 0.0720 0.0730
Burgers’ $d=2$ Relative $\ell_{2}$ error (%) PI-DeepONet 28.48 ± 4.17 28.63 ± 4.10 28.26 ± 4.38 12.33 ± 5.14 -
SepONet 27.79 ± 4.40 28.16 ± 4.24 22.78 ± 6.47 10.25 ± 4.44 7.51 ± 4.04
RMSE (×$\times 10^{-2}$) PI-DeepONet 4.09 ± 2.77 4.11 ± 2.78 4.07 ± 2.78 1.96 ± 1.92 -
SepONet 4.01 ± 2.75 4.05 ± 2.76 3.30 ± 2.55 1.65 ± 1.64 1.23 ± 1.44
Memory (GB) PI-DeepONet 5.085 9.695 17.913 35.433 -
SepONet 0.605 0.607 0.607 0.609 0.716
Training time (hours) PI-DeepONet 0.5135 1.3896 2.6904 5.923 -
SepONet 0.0725 0.0707 0.0703 0.0612 0.0605
Nonlinear diffusion $d=3$ Relative $\ell_{2}$ error (%) PI-DeepONet - - - - -
SepONet 31.94 ± 9.18 25.48 ± 8.95 21.16 ± 7.82 10.21 ± 3.31 6.44 ± 1.69
RMSE (×$\times 10^{-2}$) PI-DeepONet - - - - -
SepONet 3.44 ± 1.13 2.73 ± 0.99 2.27 ± 0.91 1.09 ± 0.32 0.68 ± 0.15
Memory (GB) PI-DeepONet - - - - -
SepONet 2.923 3.139 4.947 6.995 13.139
Training time (hours) PI-DeepONet - - - - -
SepONet 0.1175 0.1408 0.1849 0.3262 0.5575
Appendix C Complete solution to separation of variables example Eq. 7
Recall the linear PDE system treated in Section 2.3:
(46) $\mathcal{M}[t]s(y)=\mathcal{L}_{1}[y_{1}]s(y)+\cdots+\mathcal{L}_{d}[y_{d}]s(y),$
where $\mathcal{M}[t]=\frac{d}{dt}+h(t)$ is a first order differential operator of $t$, and $\mathcal{L}_{1}[y_{1}],...,\mathcal{L}_{d}[y_{d}]$ are linear second order ordinary differential operators of their respective variables $y_{1},...,y_{d}$ only. Furthermore, assume we are provided Robin boundary conditions in each variable and separable initial condition $s(t=0,y_{1},\ldots,y_{d})=\prod_{n=1}^{d}\phi_{n}(y_{n})$ for functions $\phi_{n}(y_{n})$ that satisfy the boundary conditions.
Assuming a separable solution exists, $s(y)=T(t)Y_{1}(y_{1})\cdots Y_{d}(y_{d})$, the PDE can be decomposed in the following form:
(47) $\frac{\mathcal{M}T(t)}{T(t)}=\frac{\mathcal{L}_{1}Y_{1}(y_{1})}{Y_{1}(y_{1})}+\cdots+\frac{\mathcal{L}_{d}Y_{d}(y_{d})}{Y_{d}(y_{d})},$
where it is apparent that each term in the sequence is a constant, since they are each only functions of a single variable. Consequently, we may solve each of the $\mathcal{L}_{n}$ terms independently using Sturm-Liouville theory. After we have found the associated eigenfunctions ($Y_{n}^{k_{n}}$) and eigenvalues ($\lambda_{n}^{k_{n}}$), we may manually integrate the left-hand side. Finally, we may decompose the separable initial condition into a product of sums of the orthonormal basis functions (eigenfunctions) of each variable. The resulting solution is given by
[…]
Here, $k=(k_{1},\ldots,k_{d})$, where $k_{n}\in\{1,2,...,\infty\},\forall n\in\{1,...,d\}$, is a lumped index that counts over all possible products of eigenfunctions $Y_{n}^{k_{n}}$ with associated eigenvalues $\lambda_{n}^{k_{n}}$. $\langle\cdot\rangle_{n}$ is an appropriate inner product associated with the separated Hilbert space of the $\mathcal{L}_{n}$-th operator. To obtain Eq. 7 in the main manuscript, one only need to break up the sum over all $k_{n}$ indices into a single ordered index.
Appendix D Linear Heat Equation Example
SepONet is motivated by the classical method of separation of variables, which is often employed to solve linear partial differential equations (PDEs). To illustrate the connection between these approaches, consider the linear heat equation:
(49) $\frac{\partial s}{\partial t}=\frac{1}{\pi^{2}}\frac{\partial^{2}s}{\partial x^{2}},\quad(x,t)\in(0,1)\times(0,1],$
$s(x,0)=u(x),\quad x\in(0,1),$
$s(0,t)=s(1,t)=0,\quad t\in(0,1).$
The goal is to solve this equation for various initial conditions u(x) using both the separation of variables technique and the SepONet method, allowing for an intuitive comparison between the two.
D.1 Separation of Variables Technique
We seek a solution in the form:
(50) $s(x,t)=X(x)T(t)$
for functions $X$, $T$ to be determined. Substituting Eq. 50 into Eq. 49 yields:
(51) $\frac{X^{{}^{\prime\prime}}}{X}=-\lambda~{}~{}\text{and}~{}~{}\frac{\pi^{2}T^{{}^{\prime}}}{T}=-\lambda$
for some constant $\lambda$. To satisfy the boundary condition, $X$ must solve the following eigenvalue problem:
(52) $X^{{}^{\prime\prime}}(x)+\lambda X(x)$ $=0,\quad x\in(0,1),$
$X(0)=X(1)$ $=0,$
and $T$ must solve the ODE problem:
(53) $T^{{}^{\prime}}(t)=-\frac{\lambda}{\pi^{2}}T(t).$
The eigenvalue problem Eq. 52 has a sequence of solutions:
(54) $\lambda_{k}=(k\pi)^{2},\quad X_{k}(x)=\sin(k\pi x),\quad\text{for}~{}~{}k=1,2,\dots$
For any $\lambda$, the ODE solution for $T$ is $T(t)=Ae^{-\frac{\lambda}{\pi^{2}}t}$ for some constant $A$. Thus, for each eigenfunction $X_{k}$ with corresponding eigenvalue $\lambda_{k}$, we have a solution $T_{k}$ such that the function
(55) $s_{k}(x,t)=X_{k}(x)T_{k}(t)$
will be a solution of Eq. 51. In fact, an infinite series of the form
(56) $s(x,t)=\sum_{k=1}^{\infty}X_{k}(x)T_{k}(t)=\sum_{k=1}^{\infty}A_{k}e^{-k^{2}t}\sin(k\pi x)$
will also be a solution satisfying the differential operator and boundary condition of the heat equation Eq. 49 subject to appropriate convergence assumptions of this series. Now let $s(x,0)=u(x)$, we can find coefficients:
[…]
such that Eq. 56 is the exact solution of the heat equation Eq. 49.
D.2 SepONet Method
In this section, we apply the SepONet framework to solve the linear heat equation Eq. 49 and compare the basis functions it learns with those derived from the classical separation of variables method. Recall that a SepONet, parameterized by $\boldsymbol{\theta}$, approximates the solution operator of Eq. 49 as follows:
(58) $G_{\boldsymbol{\theta}}(u)(x,t)=\sum_{k=1}^{r}\beta_{k}(u(x_{1}),u(x_{2}),\ldots,u(x_{128}))\tau_{k}(t)\zeta_{k}(x),$
where $x_{1},x_{2},\ldots,x_{128}$ are 128 equi-spaced sensors in $[0,1]$, $\beta_{k}$ is the $k$-th output of the branch net, and the basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ are the $k$-th outputs of two independent trunk nets.
Training settings
The branch and trunk networks each have a width of 5 and a depth of 50. To determine the parameters $\boldsymbol{\theta}$, we trained SepONet for 80,000 iterations, minimizing the physics loss. Specifically, we set $\lambda_{I}=20$, $\lambda_{b}=1$, $N_{f}=100$, and $N_{c}=128^{2}$ in the physics loss. The training functions (initial conditions) $\left\{u^{(i)}\right\}_{i=1}^{N_{f}}$ were generated from a Gaussian random field (GRF) $\sim\mathcal{N}\left(0,25^{2}\left(-\Delta+5^{2}I\right)^{-4}\right)$ using the Chebfun package [8], ensuring zero Dirichlet boundary conditions. Additional training settings are detailed in Appendix B.2 of the main text.
Evaluation
We evaluated the model on 100 unseen initial conditions sampled from the same GRF, using the forward Euler method to obtain reference solutions on a $128\times 128$ uniform spatio-temporal grid.
Impact of the rank $r$
Since $\tau_{k}(t)$ and $\zeta_{k}(x)$ are independent of the initial condition $u$, learning an expressive and rich set of basis functions is crucial for SepONet to generalize to unseen initial conditions. To investigate the impact of the rank $r$ on the generalization error, we trained SepONet with ranks ranging from 1 to 50. The mean RMSE between SepONet’s predictions and the reference solutions over 100 unseen test initial conditions was reported. For comparison, we also computed the mean RMSE of the truncated analytical solution at rank $r$ for $r$ from 1 to 15. The results are presented in Fig. 4.
As $r$ increases, the truncated analytical solution quickly converges to the reference solution. The nonzero RMSE arises due to numerical errors in computing the coefficients $A_{k}$ and the inherent inaccuracies of the forward Euler method used to generate the reference solution. For SepONet, we observed that when $r=1,2$, the mean RMSE aligns closely with that of the truncated solution. However, as $r$ increases beyond that point, the error decreases more gradually, stabilizing around $r=50$. This indicates that SepONet may not necessarily learn the exact same basis functions as those from the truncated analytical solution. Instead, a higher rank $r$ allows SepONet to develop its own set of basis functions, achieving similar accuracy to the truncated solution.
SepONet basis functions
The learned basis functions for different ranks $r$ are visualized in Fig. 5 to Fig. 9.
At $r=1$, SepONet learns basis functions that closely resemble the first term of the truncated solution. For $r=2$, the learned functions are quite similar to the first two terms of the truncated series. However, when $r=5$, the basis functions diverge from the truncated solution series, although some spatial components still resemble sinusoidal functions and the temporal components remain monotonic. As r increases further, SepONet continues to improve in accuracy, though the learned basis functions increasingly differ from the truncated series, confirming SepONet’s ability to accurately approximate the solution using its own learned basis functions.
Figure 4: Comparison of RMSE between the truncated analytical solution and SepONet predictions for varying rank $r$. The truncated analytical solution quickly converges, while SepONet shows a slower decay in error, converging around $r=50$.
Figure 5: Learned basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ for $r=1$. SepONet learns the same basis functions as the first term of the truncated solution.
Figure 6: Learned basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ for $r=2$. SepONet learns very similar basis functions as the first two terms of the truncated solution.
Figure 7: Learned basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ for $r=5$.
Figure 8: Learned basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ for $r=10$.
Figure 9: Learned basis functions $\tau_{k}(t)$ and $\zeta_{k}(x)$ for $r=50$.
Appendix E Visualization of SepONet Predictions
In this section, we showcase the performance of trained SepONets in predicting solutions for PDEs under previously unseen configurations. The SepONets were trained using $N_{f}=100$ and $N_{c}=128^{d}$, where $d$ denotes the dimensionality of the PDE problem. The prediction results are presented in Figs. 10, 11, 12, and 13.
Figure 10: (1+1)-$d$ Diffusion-reaction equation.
Figure 11: (1+1)-$d$ Advection equation.
Figure 12: (1+1)-$d$ Burgers’ equation.
Figure 13: (2+1)-$d$ Nonlinear diffusion equation. Two snapshots at $t=0.5$ and $t=1$ are presented.